### v1

# Testing Hypotheses of Covariate Effects on Topics of Discourse

Abstract
We introduce an approach to topic modelling with document-level covariates that remains tractable in the face of large text corpora. This is achieved by de-emphasizing the role of parameter estimation in an underlying probabilistic model, assuming instead that the data come from a fixed but unknown distribution whose statistical functionals are of interest. We propose combining a convex formulation of non-negative matrix factorization with standard regression techniques as a fast-to-compute and useful estimate of such a functional. Uncertainty quantification can then be achieved by reposing non-parametric resampling methods on top of this scheme. This is in contrast to popular topic modelling paradigms, which posit a complex and often hard-to-fit generative model of the data. We argue that the simple, non-parametric approach advocated here is faster, more interpretable, and enjoys better inferential justification than said generative models. Finally, our methods are demonstrated with an application analysing covariate effects on discourse of flavours attributed to Canadian beers.
Keywords: Text analysis, topic models, non-negative matrix factorization, bootstrapping, numerical linear algebra, text as outcomes
1 INTRODUCTION
Recent decades have witnessed a paradigm shift in statistics in which the central notion of data has expanded to include such diverse objects as images, audio recordings, and unstructured text. Not incidentally, this shift has fostered an increasing overlap between statistics and neighbouring fields. An influential product of this interaction for the analysis of text data is the field of topic modelling, where practitioners assume that a collection of documents can be reduced to a low-dimensional “topic” representation. Loosely, this can be thought of as unveiling the central themes that permeate a collection of documents. Once compressed in this way, documents can be sorted, searched, and further manipulated for downstream information processing tasks. They can also be mined for underlying regularities that provide statistical insights into the nature of the content.
The most pervasive paradigm for topic modelling involves the specification of a generative probabilistic model, referred to as a Probabilistic Topic Model (PTM). Some well-known PTMs include Probabilistic Latent Semantic Indexing (Hofmann, 1999), Latent Dirichlet Allocation (Blei et al., 2003), and Correlated Topic Models (Blei and Lafferty, 2005). A survey of these and related methods can be found in Blei (2012). All of the aforementioned methods model a document’s words as being drawn from a mixture distribution over the vocabulary of possible words. The mixture components are referred to as topics, as they constitute a small number of probability vectors representative of the entire corpus. The mixture probabilities associated with these components give the prevalence of each topic amongst the documents. By estimating both the topics and their probability from raw text data, users obtain a nuanced summary of a corpus’ content without painstakingly sifting through it manually. We are especially interested in extensions of these models that allow for covariates to influence the probability of topics within the corpus; the most prominent PTM-based framework for this purpose is the Structural Topic Model (Roberts et al., 2016; Schulze et al., 2023). This is counter to the earlier Supervised Topic Model (Blei and McAuliffe, 2007) and related models (Taddy, 2013) (Egleston et al., 2021)), which treat topics as covariates for an external response of interest. Unfortunately, the inclusion of covariate effects immediately renders such PTMs difficult to use in practice. Getting around this issue is the main focus of the paper.
The issue of estimating covariate effects on topics in documents is motivated by questions of regional differences in the discourse of flavours of Canadian craft beers. grains are predominantly grown in the prairies and the freshness of the ingredients may result in beers that have grainier, maltier flavours. While most breweries will produce a suite of beer styles including India Pale Ales, Stouts, Pilsners, etc $\ldots$, we hypothesize that the flavours within those beer categories will tend towards regional effects that can be quantified from the text used to describe the beer flavours. Previous work on single malt scotch whiskies used presence or absence of specific terms as covariates (Lapointe and Legendre, 1994), however that dataset was produced by a single reviewer based on a standardized language. Our dataset is the unstructured text from thousands of different product reviewers, making binary covariates a less useful measure compared to covariates based on the strength of topics.
The difficulties in including covariate effects manifest in two, related ways. First is the challenge of implementing bespoke approximate inference techniques, like Markov chain Monte Carlo (MCMC) and variational inference, that are needed to fit PTMs incorporating the desired extra structure. Fortunately, research in automating this step of the modelling pipeline has proliferated in the last decade; see, for example, advances in scalable variational inference (Kucukelbir et al., 2017; Ranganath et al., 2014) and their implementation in increasingly-popular probabilistic programming languages such as Stan (Carpenter et al., 2017). These general systems allow users to specify broad classes of probabilistic models and treat inference as a black-box procedure whose details are hidden from the analyst.
Despite this progress, a second level of difficulty remains largely unaddressed. Namely, the layers of approximation needed to fit STMs and related models is formidable, leading to doubts about the validity of inferences drawn from them. For example, due to issues of non-identifiability and non-conjugacy, Roberts et. al. employ a “partially collapsed variational expectation-maximization algorithm that uses a Laplace approximation to the nonconjugate portion of the model” (Roberts et al., 2016). Interestingly, the issue of inferential accuracy is present even in PTMs which do not exhibit covariate structure. Variational inference – the first method used to fit Latent Dirichlet Allocation – struggles to accurately capture higher moments, necessitating post-hoc corrections and diagnostics (Giordano et al., 2018; Yao et al., 2018). MCMC can be problematic too; Griffiths and Steyvers (2004) discuss difficulties applying Gibbs sampling to Latent Dirichlet Allocation, primarily due to the non-identifiability of the model’s latent variables. Gradient-based methods such as Hamiltonian Monte Carlo (Neal, 2011) offer no solution, as they struggle to handle the high dimensional discrete latent variables that are ubiquitous in PTMs. Ultimately, these computational challenges and the resulting nested approximations beg the question of whether simpler approaches are possible. In particular, this paper investigates the virtues of eschewing strict adherence to the PTM paradigm.
Instead of building a generative probabilistic model whose parameters encode information about the topics, our approach is to instead think of the documents and any associated covariates as being drawn from a fixed but unknown distribution. The reason for this shift in perspective is twofold. First, it affords greater flexibility in that we may compute and utilize any statistic we find useful to our analysis – not just those which estimate parameters under a parametric model. With this flexibility comes computational efficiency, as efforts can be focused on statistics for which fast, reliable algorithms exist. Second, the variability of these statistics can be estimated using non-parametric resampling techniques like the bootstrap. This philosophy of inference is illustrated by an analogy to standard OLS regression: one need not assume that the true data-generating mechanism is given by a Gaussian linear model for the OLS estimates to usefully summarize the data. Instead, one can treat the best-fit regression line as a statistic whose sampling distribution under the true, unknown model is to be estimated. For a discussion of inference in this vein, see Taddy et al. (2016).
In light of the challenges posed by PTMs, we propose casting covariate-informed topic based on non-negative matrix factorization (NMF) and regression models. NMF (Lee and Seung, 1996), (Lee and Seung, 1999; Lee and Seung, 2001) has a long history of applications in text analysis, and can be viewed as a linear algebraic analogue of Latent Dirichlet Allocation. Unfortunately, NMF requires solving a non-convex optimization problem. While this poses computational challenges, it also impedes treating the solution as a statistic – random initialization and convergence to local optima mean that the optimization procedure cannot be viewed as a deterministic mapping of the input data. Recent work by the theoretical computer science community on separable NMF addresses this problem. By imposing additional constraints on the solution, this formulation admits a convex optimization problem, thereby retaining the view of the output as a function of the data. This intermediary statistic is then passed to a regression framework, modelling the relationship between topics and document-level covariates. Passage to OLS permits inference via a fast bootstrap algorithm which would otherwise be intractable. Alternatively, normalizing the topic prevalence matrix allows inference through Beta regression. We call this combined suite of tools Bootstrapped (or Beta) Regression Effects Topic Trends (BRETT). Our software package is available for download at https://github.com/iamdavecampbell/NMFregress.
The paper is organized as follows. Section (2) establishes notation and lays out the general mathematical setup before providing a brief account of non-negative matrix factorization and its ability to find latent structure in text data. The section continues by outlining how separable NMF leads to the concept of anchor words and tractable algorithms that solve the NMF problem uniquely and globally. Section (3) leverages these developments in connecting NMF to ordinary regression models, permitting the inclusion of covariate effects (the motivating problem of this work). Statistical inference is addressed within this framework by showing how a bootstrap scheme can be added in a way that obviates repeatedly solving the NMF problem. Section (4) serves two purposes. First, BRETT is compared to the STM framework using publicly-available text data from NeurIPS conference papers. The section then applies BRETT to a data set of Canadian beer reviews, paying particular attention to regional differences as encoded by a covariate associated to each document. Conclusions and future research directions follow in section (6).
2 Methods
The term-document matrix (TDM) $\mathbf{X}\in\mathbb{N}^{V\times D}\subset\mathbb{R}_{+}^{V\times D}$ contains entries $x_{ij}$ representing the counts of word $i$ appearing in document $j$ for a corpus of $V$ unique words and $D$ documents. Note that substantial pre-processing is often required to transform a raw corpus into a form amenable to analysis. As an auxiliary piece of data, let $\mathbf{Z}\in\mathbb{R}^{D\times P}$ be a model matrix encoding document-level covariates (these can be discrete or continuous) and an intercept term. The only stochastic assumption we employ is that
$(\mathbf{X},\mathbf{Z})\sim\pi(\mathbf{X},\mathbf{Z})$
where $\pi$ is some fixed but unknown probability measure. Formally, this measure is required to be sufficiently well-behaved so as to permit unproblematic bootstrapping. For standard covariate-free topic modelling, one could compute a statistic
$\psi_{1}\left(\mathbf{X}\right)$
that mimics the point estimates generated when fitting a PTM and thus accurately reflects the kind of correlation structures associated with the term “topic modelling”. Extending this to the case with covariates,
$\psi_{2}\left(\mathbf{X},\mathbf{Z}\right),$
serves the same purpose but also incorporates $\mathbf{Z}$ in an appropriate way. Finally, estimate the sampling distribution of these statistics by bootstrapping or other means. This sec constructs specific instances of $\psi_{1}$ and $\psi_{2}$, as well as efficient methods of estimating their sampling distributions.
2.1 Non-negative Matrix Factorization
The essential feature of NMF is its attempt to factor $\mathbf{X}$ into two “simpler” matrices $\bm{\Phi}$ and $\bm{\Theta}$, also containing non-negative entries,
$\mathbf{X}\approx\bm{\Phi}\bm{\Theta}$
where $\bm{\Phi}\in\mathbb{R}_{+}^{V\times T}$, $\bm{\Theta}\in\mathbb{R}_{+}^{T\times D}$, and $T\ll V$ controls the rank of the factorization. A single document $\mathbf{x}_{i}$ can then be written as $\mathbf{x}_{i}\approx\bm{\Phi}\bm{\theta}_{i}$, or equivalently
$\mathbf{x}_{i}\approx\theta_{i1}\bm{\phi}_{1}+\theta_{i2}\bm{\phi}_{2}+\ldots+\theta_{iT}\bm{\phi}_{T}.$ (1)
NMF represents documents as non-negative combinations of a small number of representative vectors encoded in the columns of ${\bm{\Phi}}$. These representative vectors constitute low-rank structure that provide useful summaries of the original, high-dimensional data. In accordance with PTMs, the $T$ columns of ${\bm{\Phi}}$ are referred to as “topics” where a particular topic ${\bm{\phi}}_{i}$ contains normalized “pseudo-counts” or “weights” associated with each word in the vocabulary. These weights dictate the relative importance of words within each topic. From (1), ${\bm{\Theta}}$ controls how these topics are allocated throughout a document, thereby defining the relative importance of topics within documents. The approximate factorization is the solution to the following optimization problem:
$\underset{{\bm{\Phi}},{\bm{\Theta}}}{\text{argmin}}\quad{\left\lVert\mathbf{X}-{\bm{\Phi}\Theta}\right\rVert}_{\mathrm{F}}$
$\text{subject to}\quad{\bm{\Phi}},{\bm{\Theta}}\succeq 0,$ (2)
where ${\left\lVert\hskip 2.84526pt\cdot\hskip 2.84526pt\right\rVert}_{\mathrm{F}}$ is the Frobenius norm, though others norms could be used in bespoke applications. Unfortunately, (2.1) is non-convex and thus NP-hard in general (Vavasis, 2010). Typically, alternating minimization schemes are used to find a local minimum, and experience shows this does give sensible topic representations in practice. As mentioned in the introduction, we are more interested in the ability to treat (2.1) as defining a statistic, which is not possible in the non-convex regime.
2.2 Separable NMF
Treating (2.1) as defining statistics requires appealing to the concept of anchor words, central in formulating separable NMF. First studied by Donoho and Stodden (Donoho and Stodden, 2004), the separability assumption states that within ${\bm{\Phi}}$ lies a $T\times T$ diagonal matrix, possibly after permuting ${\bm{\Phi}}$’s rows appropriately. In the language of topics, this means that for each topic (column of ${\bm{\Phi}}$) there exists a word with non-zero weight only within that topic. More formally, for each column $j$ we can find an entry $\phi_{ij}>0$ such that $\phi_{kj}=0$ for all $k\neq i$. Such a word is called an anchor word, for it “ties down” a particular topic.
Anchor words have been studied and used extensively in developing provably efficient algorithms for estimating PTMs (Arora et al., 2012; Arora et al., 2018; Gillis, 2014). Formulated within the PTM framework, these algorithms can recover the underlying model’s parameters given the assumption that anchor words really exist. In contrast, we treat the presence of anchor words as an additional constraint on $\mathbf{\Phi}$ when solving (2.1). To see why this aids in tractability, suppose that $\mathbf{X}$ is such that $\mathbf{X}={\bm{\Phi}}{\bm{\Theta}}$ exactly. The existence of anchor words implies
$\mathbf{\Phi}=\left[\begin{array}[]{c}\mathbf{\Lambda}\\ \mathbf{\Gamma}\end{array}\right],\mbox{ where }\mathbf{\Lambda}=\left[\begin{array}[]{cccc}\lambda_{1}&0&\cdots&0\\ 0&\lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\lambda_{T}\end{array}\right]$
and $\mathbf{\Gamma}$ is the remaining block of $\mathbf{\Phi}$ with entries $\gamma_{ij}$. The anchor words appear in the first $T$ rows of $\bm{\Phi}$, a fact that can always be enforced with an appropriate permutation. Then
$\mathbf{X}=\left[\begin{array}[]{c}\mathbf{\Lambda}\\ \mathbf{\Gamma}\end{array}\right]\left[\begin{array}[]{c}\mathbf{\Theta}\end{array}\right]$ $=\left[\begin{array}[]{c}\mathbf{\Lambda}\mathbf{\Theta}\\ \mathbf{\Gamma}\mathbf{\Theta}\end{array}\right]$
$=\left[\begin{array}[]{c}\mathbf{X}_{\dagger}\\ \mathbf{\Gamma}\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}\end{array}\right],$
where $\mathbf{X}_{\dagger}$ is the restriction of $\mathbf{X}$ to its first $T$ rows – those corresponding to the anchor words. Separability thus means that $\bm{\Theta}$ is sitting inside $\mathbf{X}$ (up to a diagonal scaling). Defining $\mathbf{X}_{\ddagger}=\mathbf{\Gamma}\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$ as the non-anchor block of $\mathbf{X}$, the non-anchor rows are then seen to be non-negative combinations of the anchor rows.
All this assumes that the factorization holds exactly, which will never be true in practice. However, the above insights motivate adding separability as a constraint to (2.1). Requiring that $\mathbf{X}_{\dagger}=\bm{\Lambda}\bm{\Theta}$, one can focus exclusively on the non-anchor block to minimize
${\left\lVert\mathbf{X}_{\ddagger}-\mathbf{\Gamma}\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}\right\rVert}_{\mathrm{F}}.$
Setting $\mathbf{Y}=\mathbf{\Gamma}\left(\bm{\Lambda}\right)^{-1}$ highlights how the anchor block terms, $\mathbf{X}_{\dagger}$, are predictive of the within topic terms, $\mathbf{X}_{\ddagger}$,:
$\hat{\mathbf{Y}}=\underset{\mathbf{Y}}{\text{argmin}}\quad{\left\lVert\mathbf{X}_{\ddagger}-\mathbf{Y}\mathbf{X}_{\dagger}\right\rVert}_{\mathrm{F}}$
$\text{subject to}\quad\hat{\mathbf{Y}}\succeq 0.$ (3)
Solving (2.2) amounts to a series of non-negative least squares (NNLS) problems, one for each of the $V-T$ non-anchor rows in $\mathbf{X}$. NNLS is well-studied and – crucially – convex (Chen and Plemmons, 2009); this establishes the overall convexity of (2.1) with the additional anchor word constraint. As (2.2) attempts to write the non-anchor words as non-negative combinations of the anchors, it follows that anchor words determine the semantic properties of each topic.
Having found a globally optimal $\hat{\mathbf{Y}}$, one final constraint must be added to uniquely determine $\bm{\Phi}$ and $\bm{\Theta}$. Forcing $\bm{\Phi}$ to have columns summing to 1, giving it the interpretation of a matrix of probabilities of terms within topics, write
${\left\lVert\bm{\phi}_{j}\right\rVert}_{1}=\lambda_{j}+\sum_{i=T+1}^{V}\gamma_{ij}=1,$
whence we find
$\lambda_{j}=1-\sum_{i=T+1}^{V}\gamma_{ij}=1-\lambda_{j}\sum_{i=T+1}^{V}\left(\frac{\gamma_{ij}}{\lambda_{j}}\right)$
and notice the $(i,j)^{\text{th}}$ element of $\mathbf{\Gamma}\left(\mathbf{\Lambda}\right)^{-1}=\mathbf{Y}$ in the summation. This gives
$\lambda_{j}=\frac{1}{1+\sum_{i=T+1}^{V}\hat{y}_{ij}}.$
Besides permitting unique determination of $\bm{\Lambda}$, the sum-to-one constraint also means the columns of $\bm{\Phi}$ can be interpreted probabilistically, making comparisons to conventional PTMs easier. Given anchor words, we solve for $\hat{\mathbf{Y}}$, followed by $\hat{\bm{\Lambda}}$, $\hat{\bm{\Gamma}}$, $\hat{\bm{\Phi}}$, and $\hat{\bm{\Theta}}$ in turn as outlined in Algorithm 1.
[…]
2.3 Determining Anchor Words
The above discussion assumes anchor words are known in advance. In practice, they must be found or – as is done here – choosen.
In some circumstances the analyst may have strong interests in using hand select anchors for subsequent inference. In our experience, a better approach is to employ a data driven tool for procuring anchors. Topics represent a set of potential features from which the analyst can choose as interpretable, inferential measures. A data driven approach also provides the analyst with insight into the appropriateness of the text for exploring hypotheses of interest.
Unlike in the application of anchor words to PTMs, there is no probabilistic model in which to posit their existence. Instead, the Successive Projection Algorithm (SPA) is used to find a set of $T$ words to serve as anchors (Gillis, 2014; Bezerra et al., 2001). SPA and variants thereof are widely used in the probabilistic setting (Arora et al., 2012; Arora et al., 2018), but here the algorithm is treated as a method of variable selection. As such, it is a pre-processing step that must be undertaken before solving (2.1). All further analysis, including resampling-based inference, is then conditional on these anchor words. While this may seem unmotivated, note that it is actually crucial to interpretability. Elaboration of this point is deferred to section (3).
SPA can be informally described by considering the $V$ rows of $\mathbf{X}$ as vectors in $\mathbb{R}^{D}$.
  • Take the first anchor to be the point farthest from the origin; by the definition of the 2-norm, this is just the word that is most often within documents in the corpus. Note that non-informative “stopwords” are removed as a pre-processing step prior to constructing $\mathbf{X}$.
  • The next anchor is the word farthest from the line spanned by the first.
  • The third is the word farthest from the plane spanned by the first two.
  • Repeat until $T$ anchor words are found.
While in the form above SPA only relies on standard linear algebra calculations and is equivalent to the pivoting strategy employed by many QR-decomposition algorithms (Engler, 1997; Businger and Golub, 1965; Gillis, 2014). Computing the QR decomposition of $\mathbf{X}^{\mathrm{T}}$ and extracting the pivots provides a permutation on the set $\{1,2,\ldots,V\}$ whose first $T$ elements give the row indices of our chosen anchor words in $\mathbf{X}$.
While SPA was used in the probabilistic regime because (under suitable regularity conditions) it provably uncovers the “true” anchor words, viewing it in this light reveals why it is useful without an underlying model. As noted by Businger and Golub (1965), column-pivoted QR (and by extension SPA) finds rows of $\mathbf{X}$ that are “very linearly independent.” Another useful perspective is that SPA finds words which trade off relevance (i.e., words that are used frequently in the corpus) with semantic dissimilarity (words that make a large angle with the subspace spanned by the previously chosen anchors). Much work has been done on extending column-pivoted QR for variable selection and giving bounds on its success (Gu and Eisenstat, 1996; Chan, 1987; Broadbent et al., 2010), but the basic version is known to work well in practice (Gillis, 2014). Ultimately, the QR connection affords a computational advantage in that fast implementations, like that of LAPACK (Anderson et al., 1999), can be used for off-the-shelf anchor selection.
The NMF-based topic modelling approach is summarized in algorithm (1). All subroutines can be reliably computed in polynomial time using established linear algebra and optimization packages. While it is not in itself the main purpose of the paper, algorithm (1) should be of interest to practitioners interested in fast, reproducible alternatives to traditional topic models. We caution such readers not to analogize too closely with PTMs, as this can lead to confusion. Because the anclor words are defined as being predictive of the rest of the TDM, the presence of anchor words completely determines whether a topic is “present” in a given document; this follows from $\bm{\Lambda}$’s diagonal form. Thus, documents without any anchor words correspond to a column of $\bm{\Theta}$ with no non-zero entries: the document contains “no topics.” This can be puzzling if one is wedded to the PTM perspective; indeed, such behaviour is impossible in such models. However here, such documents contain no relevant information for the topic. This could be a genuine problem in corpora with very small documents, but experience shows that in reasonably sized collections (such as the ones we study later) these “empty” documents are rare. Alternatively, the raw vocabulary could be transformed into BERT word embeddings Devlin et al. (2019) before constructing the TDM Dieng et al. (2020).
Input: a TDM $\mathbf{X}$.
  • Compute the QR decomposition of $\mathbf{X}^{\mathrm{T}}$ (with column pivoting).
  • Extract the anchor indices $\{a_{1},a_{2},\ldots,a_{T}\}\subset\{1,2,\ldots,V\}$ from the pivoting strategy.
[…]
  • For $j\in\{1,2,\ldots,T\}$, compute $\hat{\lambda}_{j}=1\big{/}\left(1+\sum_{i=T+1}^{V}\hat{y}_{ij}^{\text{opt.}}\right)$.
  • Compute:
    • $\circ$
      $\hat{\bm{\Lambda}}=\mathrm{diag}(\hat{\lambda}_{1},\ldots,\hat{\lambda}_{T})$.
  • $\circ$
    $\hat{\bm{\Gamma}}=\hat{\mathbf{Y}}^{\text{opt.}}\bm{\hat{\Lambda}}$.
  • $\circ$
    $\hat{\bm{\Theta}}=\left(\hat{\bm{\Lambda}}\right)^{-1}\mathbf{X}_{\dagger}$.
  • $\circ$
    $\bm{\Phi}$, by concatenating $\bm{\Lambda}$ and $\bm{\Gamma}$ column-wise.
    Output: $\psi_{1}=\left\{\hat{\bm{\Phi}},\hat{\bm{\Theta}}\right\}$.
  • Algorithm 1 NMF-based topic modelling with anchor words
    2.3.1 Selecting the Number of Topics
    In unsupervised methods like clustering and topic modelling, there is no ground truth with which to compare. Several measures have been proposed to assess the number and quality of topics. Perplexity measures how well a held out set of documents conform to the topic model, however it has been shown to correlate poorly with human interpretable topics (Chang et al., 2009). UMass coherence considers the frequency of occurence of within-topic word pairs as a measure of the unique interpretability of the topic construction as assessed by the log ratio of the number documents in which the most important words co-occur compared to number of documents containing a word (Mimno et al., 2011). Variations attempt to improve the measure’s ability to capture what humans consider to be good topics (Newman et al., 2010; Röder et al., 2015). However, it is notable that the target of these methods remains the ability for a human to interpret topics and consider them to be appropriate as compared to domain expertise.
    For the purposes of inference, the topics represent the metric by which a hypothesized effect can be measured. As with any quantitative discipline, the scientist must determine if the metric is appropriate and targets, at least in proxy, the analytic goal.
    Well curated data that is specific to the analytic goal should be combined with a moderate number of topics relative to the number of documents. This is explored empirically in Section 5.4.
    3 Modelling Associations Between Topics and Covariates
    Matrix $\bm{\Phi}$ defines the topics, while $\bm{\Theta}$ defines the extent to which a document uses a topic. Modelling focuses on assessing the extent to which covariates affect $\bm{\Theta}$. This section outlines a fast OLS routine for the special case of categorical covariates and then continues into more realistic regression models for continuous covariates.
    When cast as a PTM, categorical covariates in an Ordinary Least Squares (OLS) hold the interpretation that coefficients are the estimated deviation of the mean in moving in or out of a category. Maintaining this interpretability through confidence intervals necessitates the use of bootstrap confidence intervals outlined in Section 3.1. This is expanded into the general purpose Beta-regression approach outlined in Section 3.2.
    3.1 OLS Regression and Bootstrapping
    Focusing on $\bm{\Theta}$ (as extracted by algorithm (1)) with elements $\theta_{ij}$ describing the probability of topic $i$ in document $j$ and covariate vector $\mathbf{z}^{j}$ is the $j^{\text{th}}$ row of $\mathbf{Z}$, write
    $\widetilde{\theta}_{ij}\approx\langle\mathbf{z}^{j},\bm{\beta}_{i}\rangle,$
    where $\bm{\beta}_{i}\in\mathbb{R}^{P}$ is the coefficient vector of interest, and $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{j}\theta_{ij}$. It can be shown that this normalization scales $\bm{\beta}_{i}$ by a constant, which has downstream computational benefits while also ensuring its interpretation as a matrix of probabilities of topics within documents. Formalizing the above, the goal is to solve for $\mathbf{B}$ with columns $\beta_{i}$ by solving $T$ OLS problems,
    $\text{minimize}\quad{\left\lVert\widetilde{\bm{\Theta}}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$ (4)
    The constraint $10<\theta_{ij}<$ 1 imposes heteroskedasticity and directionality to the residuals of (4). The breach of the Gauss-Markov assumptions degrades the quality of OLS interval estimates. Instead bootstrap intervals are used.
    The decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$ where $\left(\bm{\Lambda}\right)^{-1}$ is diagonal scales the rows of $\mathbf{X}_{\dagger}$, an effect that is undone by row normalization. Thus, (4) can be reformulated as
    $\text{minimize}\quad{\left\lVert\widetilde{\mathbf{X}}_{\dagger}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$
    Crucially, this only involves the scaled TDM. The upshot is that algorithm (1) need not be re-run when using resampling methods to estimate the sampling distribution of $\mathbf{B}$ – nowhere does it depend on $\bm{\Theta}$.
    Algorithm (2) describes a bootstrapping procedure that takes advantage of this simplification. Resampling the documents, columns of $\mathbf{X}$, allows efficient estimation of the sampling distribution of $\psi_{2}$.
    Input: Observed TDM $\mathbf{X}$, number of bootstrap samples $b$, anchor indices $\{a_{1},a_{2},\ldots,a_{T}\}$ (as found by algorithm (1) or some other method).
    • For $i\in\{1,2,\ldots,b\}$:
      • $\circ$
        Sample $S_{1},S_{2},\ldots,S_{D}\sim\{1,2,\ldots,D\}$.
      • $\circ$
        Form $\mathbf{X}^{S}$ by concatenating $\mathbf{x}_{S_{1}},\mathbf{x}_{S_{2}},\ldots,\mathbf{x}_{S_{D}}$ row-wise.
      • $\circ$
        Form $\mathbf{Z}^{S}$ by similarly concatenating $\mathbf{z}_{S_{1}},\mathbf{z}_{S_{2}},\ldots,\mathbf{z}_{S_{D}}$.
      • $\circ$
        Form $\widetilde{\mathbf{X}}_{\dagger}^{S}$ analogous to $\widetilde{\mathbf{X}}_{\dagger}$, using $\mathbf{X}^{S}$ in place of $\mathbf{X}$. Recall from algorithm (1) that $\mathbf{X}_{\dagger}$ is formed by restricting $\mathbf{X}$ to the anchor rows $\{a_{1},a_{2},\ldots,a_{T}\}$.
      • $\circ$
        Find $\mathbf{B}^{i}=\arg\underset{\mathbf{B}}{\min}\hskip 5.69054pt{\left\lVert\widetilde{\mathbf{X}}_{\dagger}^{S}-\mathbf{B}^{\text{T}}\left(\mathbf{Z}^{S}\right)^{\text{T}}\right\rVert}_{\mathrm{F}}$ by solving $T$ OLS problems.
    Output: $\{\psi_{2}^{i}=\mathbf{B}^{i}\}_{i=1}^{B}$.
    Algorithm 2 BRETT: Bootstrapped/Beta Regression Effects Topic Trends.
    Algorithm (2) takes as input the anchor rows $\{a_{1},a_{2},\ldots,a_{T}\}$ and produces bootstrap estimates of $\mathbf{B}$’s sampling distribution conditional on the anchor words. This is an important step to sensible inference; note that without fixing the anchor words, the semantics of the various topics would change with each bootstrap iterate. Since new anchor words could be produced with each bootstrap sample (and anchor words determine the semantic coherence of a topic), the meaning of the topic indexed by $i$, say, would be in constant flux. While we find that using SPA to pre-select anchor words works well in practice, it could be altered with a bespoke tool without breaking the essential inferential mechanism. So long as “good” anchor words are produced (in the sense that the analyst finds them an informative summary of their corpus), useful for hypothesis testing and effect estimation can be devised without re-computing anchor words at each bootstrap iterate.
    3.2 Beta Regression
    Again focusing on attention on $\bm{\Theta}$ (as extracted by algorithm (1)), recall that the entries $\theta_{ij}\in[0,1]$ describes the probability (or “weight”) assigned to topic $i$ in document $j$. After selecting topic $i$ of interest to the analyst, and normalizing $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{i}\theta_{ij}$ so that rows sum to 1, vector $\widetilde{\theta}_{i}$ contains elements interpretable as the probability a word from document $j$ as being from topic $i$. For a fixed $i$, modelling $\theta_{ij}$ as depending on a vector of covariates $\mathbf{z}^{j}$ through coefficient vector $\bm{\beta}_{i}\in\mathbb{R}^{P}$,
    $\widetilde{\theta}_{ij}\sim Beta(\mbox{mean}=\mu,\mbox{precision}=\sigma),with$
    $g_{\mu}(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$
    The mean-precision parameterization of the Beta provides a more intuitive interpretation of regression (Vasconcellos and Cribari-Neto, 2005). The link functions, $g_{\mu}(\mu)$ and $g_{\sigma}(\sigma)$, provide mappings from potentially different subsets of $\mathbf{z}^{j}$ into the appropriate spaces with covariate driven variation (Simas et al., 2010).
    Beta regression can be applied directly to this problem conditional on the extraction of $\bm{\Theta}$. Asymptotic standard error estimates can be obtained from the likelihood directly or through bootstrap.
    4 Experiments and Applications
    This section presents case studies showcasing BRETT. The first analysis considers the time evolution of discourse in papers from the NeurIPS conference to compare STM and BRETT. BRETT is shown to be much faster than the STM framework while retaining inferential accuracy. The second analysis considers the regional differences in vocabulary used to describe beer flavours.
    4.1 Comparisons with STMs
    The STM class of probabilistic models is that which is closest in spirit to BRETT. However, that STMs permit covariates to influence both topic content and topic prevalence. In BRETT-style terminology, both $\bm{\Theta}$ and $\bm{\Phi}$ are thought to depend on the specified design matrix. One could compute regression-style statistics using $\bm{\Phi}$ (akin to what was accomplished in the previous section) to mimic this ability, but the efficient bootstrap-based error bars would be lost, as one would need to re-compute $\bm{\Phi}$ with each bootstrap sample. Accordingly, we only compare to those aspects of STMs most directly comparable to BRETT: their topic discovery and incorporation of “$\bm{\Theta}$-influencing” covariates.
    Papers from the NeurIPS conference between the years 1987 and 2015 are used as a toy example. The data was obtained from the UCI Machine Learning Repository (Hamner, 2017) with light pre-processing (removal of stopwords, keeping only terms appearing 50 or more times, etc.) that is essential to any real-world topic modelling implementation. The TDM in this case consists of $V=11463$ words $D=5812$ documents. As a covariate, we use the year in which each conference proceedings was published binned into five-year intervals.
    4.1.1 Topic Extraction
    Fixing the number of topics $T=100$, tables (6) and (7) show each method’s ability to extract topics from the data. We make no claim as to 100 being the “correct” number of topics, but this choice appears to give sensible results in practice. Similarly, we elect not to compute any metrics measuring the quality of these topics. Rather, we are content to say that both methods produce what look to be very semantically coherent topics and are thus defensible methods for vanilla topic modelling tasks. That said, BRETT does have several interpretability advantages. First, anchor words provide “labels” of each topic, though this point seems moot in this example: the topics found by both methods appear easily understandable to humans. More importantly, BRETT ranks topics using the $\lambda$-criterion described in section (LABEL:sec:nmf), a feature lacking in PTMs that emphasizes how predictive the anchor is of its constituent words. BRETT ranks specific, semantically interesting topics higher.
    Although the topics differ outside of their leading words, this ensures that they are at least somewhat semantically comparable; in the case of the “Gaussian” topic, tables (6) and (7) in appendix A suggest that both methods are referring primarily to Gaussian processes.
    Data fits are shown in Figure 1. The STM results are from OLS and Beta regression as estimated by the mean of 10,000 draws from the variational posterior and computing the associated regression coefficients each time. The BRETT results are from fitting Beta regression directly and from the mean of 10,000 bootstrap OLS samples. Results are similar owing to the discrete nature of the covariates.
    Uncertaintly characterization differs in interpretation across the methods. STM tools and BRETT OLS sample the regression surface producing uncertainty intervals for the model fit. BRETT Beta instead fits a Beta distribution to the data, naturally producing prediction intervals for new observations. This difference characterizes the differing widths of intervals overlayed in Figure 1.
    (a) topic ”gaussian”
    (b) Topic ”algorithm”
    Figure 1: Observed topic probabilities (orange boxplots) overlayed with boxplots of the regression surface samples or predictive interval estimates. Top row: STM, bottom row: BRETT. Columns for each subfigure are OLS (left) and Beta (right) regressions
    BRETT produces much tighter distributional estimates because STM confounds words-within-topic efects with topics-within-documents, whereas BRETT assumes a common language within topics across covariates. The difference in scale between the coefficients reflects the two algorithms’ differing fitting procedures. Both methods use sum-to-zero contrasts in constraining the single, discrete covariate, which is then interpreted as deviation in topic prevalence associated with the covariate.
    4.1.2 Computational Speed
    Using the NeurIPS data set, STMs and BRETT were compared for computational speed using the following experimental setup. The number of topics $T$ was chosen to range over the values $\{10,50,100,200\}$ and a binary variable $C\in\{T,F\}$ was devised, representing whether covariates should be included in the analysis or not. If not, classical covariate-free topic modelling was performed. Otherwise, covariate effects were estimated and 1000 simulations computed to assess the uncertainty in these estimates. In the case of BRETT, this means 1000 bootstrap samples were drawn. For the STM, “simulations” refers to 1000 draws from the posterior, which are then passed to the package’s estimateEffect(...) function. This function fits a linear model between the topic proportions (as drawn from the posterior) and covariates. STMs can be fit using several initialization schemes; we use the recommended default. Interestingly, this fits an anchor word-based PTM as a starting value for the variational inference scheme. All $4\times 2=10$ unique settings of $T$ and $C$ were timed using the same software on the same machine; figure (2) shows the results. BRETT is typically an order of magnitude faster and incurs minimal slowdowns when incorporating the extra regression functionality.
    Figure 2: Running times for BRETT and STMs with covariates (left) and without (right).
    4.1.3 Inferential Accuracy
    Beta regression is known to perform well for estimating covariate effects (Simas et al., 2010), consequently this simulation study focuses on the sampling variability of the NMF decomposition and it’s propagation through the regression procedure. Because permutations of the topic indices leave the posterior of a PTM invariant, there is no simple way to find the appropriate mapping between true and estimated parameters leaving direct comparison across iterations and between STM and NMF based regression untenable. Further complicating identifiability, rotation of the topic subspace allows topics to mix and blend into one another while the PTM maintains predicted text outcomes intact. Despite these challenges, we attempt to showcase the accuracy of BRETT using simulated data. The experimental setup involves sampling TDMs with a fixed vocabulary of size $V=1000$, altering the number of documents, $D\in\{100,500,1000\}$, where each document contains $N_{D}\in\{1000,5000,10000,15000,25000\}$ words per document. Documents were constructed with 4 topics. Topic allocation for a document comes from a Beta distribution and a linear dependence on a single, real-valued covariate transformed through a logistic link function.
    The simulation study is performed by drawing 1000 TDMs by sampling words based on the word within topic probabilities and the topic allocations based on the continuous covariate. To reduce lack of identifiability, the anchors are fixed to those used to simulate the TDMs. Given the simulation mechanism, this study explores the variability with respect to changes in $N_{D}$, and $D$ with fixed anchors, using both recalculated $\Phi$ specific to each simulation or fixing $\Phi$ at the value used to simulate the TMDs.
    To ensure a consistent simulation setup, the simulation details are somewhat atypical. First the TDM is sparsely populated, moving through the list of words and picking a random document for each in which to place a single word count. This ensures the TDMs retain the same dimension at each iteration by forcing every word to appear in at least one document under all simulations. Ensuring that all words appear at least once in the corpus alters the original probabilities of both topics within documents and terms within topics. Consequently, regression coefficients estimated from this simulation study mechanism will not be directly comparable to the ground truth used to construct the TDM. To assess the stability of the regression coefficient estimates, a pseudo-ground truth must be devised. A pseudo ground truth is constructed by summing the 1000 TDMs elementwise into the equivalent of a TMD where each document has $1000\times N_{D}$ words per document.
    The alternatives were to elimiate low probability words resulting in a much smaller vocabulary or allow a varying vocabulary size across simulations. These alternative simulation designs are also expected to alter the covariate effects from their simulated targets. The Mean Squared Error for the covariate effect effect is shown in figure 3.
    Figure 3: Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample (left) or using a single fixed $\Phi$ for all simulations (right). Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample.
    5 Application: The “Beer Data”
    A dataset about Canadian beer reviews is detailed in section 5.1 is used to show the method is able to extract well known results about flavours that define different beer styles in section 5.2. Section 5.3 tests hypotheses about whether local production of key ingredients has an affect on discourse around beer flavours. The section concludes with an examination of the stability of results under different scenarios.
    5.1 Data Acquisition and Cleaning
    Online beer review were acquired from a variety of sources respecting the sites web scraping policies as laid out in their terms and conditions and robots.txt file at the time of data collection. Data collected includes beer reviews, beer style, and the location of the brewery. After filtering to just Canadian breweries, the derived TDM consists of $V=11308$ words and $D=5168$ documents. The documents here are the concatenation of separate reviews of each beer. All text was converted to lowercase. Key n-grams such as India Pale Ale were merged into a corresponding term india_pale_ale and expanding common acronyms such as “ddh” into double_dry_hopped. Mentions of each beer’s name, style, brewery, and brewery location were deleted from the reviews as were mentions of all provinces, territories, and Statistics Canada’s Census metropolitan areas [^1] https://www.statcan.gc.ca/en/subjects/standard/daily/5000076 to avoid influencing covariate effects. Common stop words such as “the”, “of”, “and”, as well as numbers and punctuation other than those used to join n-grams was removed.
    5.2 Beer Style Hypotheses
    The dataset was filtered down to 7 beer styles, counts of which are given in table 1. The NMF model was run with 50 topics.
    The ability to test hypotheses depends on having a good metric for the characteristic of interest. The 50 topics, were examined manually to find measures of flavour attributes that coincide with the characteristics of the beer style definitions (Palmer, 2017). Some flavour attributes were spread into multiple topics, similar to the way lab sciences often indirectly measure attributes of interest from different perspectives and experiments.
    Imperial IPA IPA Lager Porter Pumpkin Ale Saison Stout
    173 603 353 188 44 257 174
    Table 1: Counts of each beer style used in Section 5.2.
    Eight hypotheses were devised based on availability of suitable metrics (topics) capturing aspects of major beer styles (Palmer, 2017).
    • IPAs and imperial IPAs are hoppy, often with tropical (H1), grapefruit (H2), or pine (H3) flavours.
    • Stouts and porters have roasted (H4), coffee (H5), and chocolate (H6) malt flavours and are often creamy (H7) in texture and mouthfeel.
    • Saisons are have citrus and peppery tones with considerable flavour influenced from the yeast (H8).
    The main terms composing the topics to test hypotheses are given in table 2. For the sake of brievity the table shows the most relevant terms from the top 20 most important terms per topic.
    The hypotheses are tested through logistic link function in Beta regression. Regression coefficients for hypotheses (H1-H8) are shown in table 3 as the deviation from the baseline beer style. In all cases our choice of baseline has a strong topic effect. Deviations from the baseline are substantial and negative except in ways that coincide with expected definitions and closely related beer styles. Imperial IPAs are considerably stronger than IPAs and often darker and with more hops. However the flavour profiles have considerable overlap compared to the within style variability in topics tropical (H1 p-value 0.08), grapefruit (H2 p-value 0.34)), and pine (H3 p-value 0.94). Porters and stouts are both made from dark roasted and chocolate malts resulting in substantial overlap in flavour compared to variability within styles leading to no significant difference in roasted (H4 p-value 0.25) or chocolate (H6 p-value 0.88) topics. Coffee complements the rich malty flavours of dark ales, but is more commonly added to stouts than in porters, resulting in a small but significant difference between beer styles in coffee (H5, p-value 0.0037). The dark malt base provides a creaminess, often enhanced by oatmeal (in stouts in particular), or the addition of nitrogen carbonation. Pumpkin ales also have creaminess from their pie-like flavours appearing as similarity to baselines (H7 difference between stouts and porters p-value .26, and difference between stouts and pumpkin ales p-value 0.070). Saisons are considerably yeastier in flavour than other beers, the closest being lagers and pumpking ales (H8 difference between saisons and lagers p-value 2e-36, and difference between saisons and pumpkin ales p-value 8e-10). Lagers are generally light in flavour and sometimes the yeast is a dominant flavour, but far from the prominance of a saison. Pumpkin ales have considerable variability spanning and could be modified from anything from a saison to a stout base. These effects are all reflected in the prediction intervals of figure 4 for topics in table 2. Beta Distributions have full support over $[0,1]$ leaving effect size prediction intervals wide with some strong prediction confidence close to 0, even when an overall effect is large as is seen in all the baseline beer styles for the hypotheses.
    topic defining terms
    grapefruit pineapple, fresh, crisp, love, citrusy, backbone, excellent, delicious, juicy, lacing, resiny, clean
    tropical mango, pineapple, fruits, lemon, gold, tangerine, dank, juicy, rind, citra, peach, papaya
    pine piney, resin, fresh, hoppy, balance, oily, west coast, balanced, huge, sticky, plenty, dipa
    yeast spice, white, pepper, lemon, spicy, wheat, golden, belgian, yellow, yeasty, banana, fruity
    brett funk, tart, funky, oak, sour, wine, tartness, lemon, apple sourness, barnyard, wild
    pumpkin cinnamon, spices, nutmeg, spice, pie, ginger, clove, amber, allspice, copper, fall, flavour
    roasted black, burnt, roasty, oatmeal, tan, thick, smooth, oats, sweetness, opaque, brown, espresso
    coffee espresso, strong, beans, brown, mouthfeel, opaque, drinkability, bodied, heavy, hints, retention, barley
    chocolate cocoa, milk, mocha, hint, black, smooth, mouthfeel, roast, characters, delicious, lacing, chocolatey
    creamy smooth, thick, brew, excellent, rich, cream, bodied, dense, sweetness, nitro, silky, feel
    Table 2: Topics used to test flavour attribute hypotheses for different beer styles. The defining words were chosen from the top 20 most predictive terms within topics.
    Hypothesis Imperial IPA IPA Lager Porter Pumpkin Ale Saison Stout
    H1 - tropical 0.28 NS -2.77$\star$ -3.07 -4.19 -4.90 -1.39 -3.75
    H2 - grapefruit -0.15 NS -2.75$\star$ -2.57 -4.94 -4.27 -2.20 -8.39
    H3 - pine 0.01 NS -3.35$\star$ -2.56 -4.52 -5.48 -2.06 -4.41
    H4 - roasted -5.10 -4.10 -3.95 -0.18NS -2.36 -3.46 -2.08$\star$
    H5 - coffee -6.17 -4.53 -5.05 -0.42 -3.83 -4.93 -2.48$\star$
    H6 - chocolate -6.32 -4.84 -6.77 -0.02NS -3.03 -4.89 -2.83$\star$
    H7 - creamy -1.05 -0.92 -1.59 -0.22NS -0.62NS -1.38 -3.36$\star$
    H8 - yeast -2.21 -2.00 -1.95 -2.68 -2.26 -1.35$\star$ -2.88
    Table 3: Beta regression coefficients for hypotheses 1-8. The baseline intercept is labelled with $\star$. Coefficients represent deviations from the baseline. Coefficients that are not significantly different from 0 at the 5% level are labelled with NS.
    Figure 4: Prediction interval for flavour topic probability for different beer styles using 50 topics.
    5.3 Regional Differences in Flavour Characteristics of IPAs
    Interest lies in assessing if there are effects of ingredient production within the flavour category. Specifically we are interested in knowing if the Prairies have grainier, breadier IPAs or if BC and Ontario have hoppier IPAs owing to localized production of base ingredients. Hypothesis tests for regional effects on flavour discourse will be confounded with differences in the types and numbers of beers produced. Focusing within a single product category, IPAs are used to test hypotheses that there are regional differences in flavour discourse. The NMF model is refitted using 776 IPAs. Using too general a corpus of all beer styles to devise topics results in messier topics that are less interpretable or useful (see for example Tshitoyan et al. (2019), Luca et al. (2022)). NMF was fitted with 50 topics resulting in topics that interpretably showcase the flavours of interest for these hypotheses.
    The grainy, bready flavour set can be assessed using earthy (H9) and caramel (H10) topics while the hoppier characteristics were assessed using grapefruit (H11) and tropical (H12). Table 4 provides characteristic terms within these topics.
    The probability of topic $Y_{i}$ for the $i^{th}$ beer is modelled with an intercept effect for the region of interest with other covariates representing the $j^{th}$ regional deviation from the intercept.
    Regression coefficients for hypotheses (H9-H12) are shown in table 4. The Prairies show significantly more earthy and caramel flavour discourse than any other region. These topics represent flavours that are not typical of the IPA style, suggesting an influence of locally produced ingredients. The grapefruit and tropical topics describe different dimensions of hop flavours characteristic of IPAs. The vast majority of beers are brewed with dried, vacuum sealed hops that travel well with reasonable shelf life. It should be expected that discourse around all IPAs mention these topics with considerable probability. It is not surprising that there are almost no significant regional differences from the BC baseline. The exception is that Ontario is showing a somewhat positive difference from BC (H11 difference between BC and Ontario p-value 0.019) in discourse of the grapefruit topic characterized by crisp, citrusy, resiny terms. Prediction intervals for effects are shown in figure 5.
    5.4 Stability With Respect to Number of Topics
    When the procedure in Section 5.2 was performed using 150 topics, the results were largely the same since additional anchors extend the topics into a higher dimensional orthogonal space. The utility of new topics starts to break down when too many topics are added because each new topic chips away at existing topics. Eventually a topic may split into two, diffusing the covariate of interest and rendering the hypothesis more challenging to define. Note that the probability of a topic for a document is monotone decreasing in the number of topics meaning that results are conditional on the model structure.
    5.5 Stability With Respect to General vs Curated Corpus
    When the topics in Section 5.3 were formed using the entire beer dataset, the topics were tuned towards separating out the broad differences, mostly driven by different beer styles. When using those topics and subsetting the documents to just IPAs, the flavour effects of interest were too diffuse to be usefully measured. Language models have been shows to have the best success when trained on data tuned to the task of interest (Tshitoyan et al., 2019).
    topic defining terms
    earthy colour, bready, leafy, grassy, grainy, smooth, lingering, english, flavours, enjoyable
    caramel malty, dark, deep, toffee, heavy, rich, buttery, bread, burnt, toasty
    grapefruit crisp, citrusy, resiny, backbone, delicious, juicy, lacing, bodied, excellent, peel
    tropical mango, dank, yellow, rind, tangerine, papaya, citra, juicy, resin, golden
    Table 4: Topics used to test flavour attribute hypotheses H9-H12 by region within the IPA beer style.
    Hypothesis BC Prairie Ontario Quebec Atlantic
    H9 - earthy -0.59 -2.72 $\star$ -0.82 -1.11 -1.46
    H10 - caramel -0.53 -3.17 $\star$ -1.01 -0.89 -0.83
    H11 - grapefruit -3.68 $\star$ 0.05 NS 0.41 -0.04 NS -0.06 NS
    H12 - tropical -3.07 $\star$ -0.19NS 0.08 NS -0.12 NS 0.33 NS
    Table 5: Beta regression coefficients for hypotheses H9-H12. The baseline intercept is labelled with $\star$. Coefficients represent deviations from the baseline. Coefficients that are not significantly different from 0 at the 5% level are labelled with NS.
    Figure 5: Prediction interval for flavour topic probability for different regions using 50 topics.
    6 Conclusion
    In natural sciences, measurements are selected based on available equipment and often hypotheses must be assessed using observable proxies of interest. Statistician are given data and analysis performed in accordance with the experimental design. In data science, the division between the data collector and data analyst is blurred. The data scientist is tasked with incorporating domain expertise into the pre-processing of the dataset. Data curation and cleaning decisions require care and will impact results. When exploring regional differences in beers, all mentions of geography had to be removed from the text to prevent mention of Ontario in the text as being the driver for the Ontario covariate effect. With text especially, data decisions impact the outcome similar to how tuning experimental equipment in lab sciences must be conditioned upon when analyzing data.
    The selection of topics as dependent variables requires care in how hypotheses are crafted to avoid the perils of using the same dataset for hypothesis generation and testing (Egami et al., 2022). The Beer hypotheses were constructed from external insights and interests around flavours, however determining anchor based topics that measure attributes of interest requires examination of $\Phi$, which in turn is based on the data. Reproducing the tests on a new independent dataset could be performed by bringing $\Phi$ or by carrying forward anchors. Even then, terminological drift suggests that application of the topic metric requires consideration of the utility of $\Phi$ within the new scientific context.
    Increased discourse about a flavour topic may or may not represent actual difference in flavour probability. Unstructured text should not be used as a replacement for scientific measurement, but should be considered as an association to be used along side domain expertise to refine hypotheses for more rigorous testing. As much as it is important to use data filtered to the context of interest to avoid introducing noisy topics, additional strength of results will be provided by refining the population of writers who provide the text data. More recent advances in Large Language Models have shown promise in categorizing text, however PTMs provide a magnitude of an observable feature rather than a binary categorical presence or absence. However there is considerable potential for exploiting LLMs as part of the data pre-processing and cleaning pipelines.
    The main advantage to the BRETT approach is its stability. The NMF decomposition with anchor words is not subject to random seeds. Anchor words can be found automatically or (partially) provided by the analyst to fine tune hypotheses. The stability in BRETT allows covariate effects to be directly modelled and uses fewer approximations than STMs.
    Appendix A PTM and STM topics
    anchor 1 2 3 4 5 6 7 8
    algorithm algorithm algorithms step online regret setting let iteration
    bound bound bounds theorem upper lower regret let lemma
    class class classes classification classifier label examples labels classifiers
    clustering clustering cluster clusters means spectral partition distance similarity
    data data analysis supervised sets test methods unlabeled missing
    distribution distribution distributions probability sample sampling prior random samples
    error error generalization errors approximation rate estimation test sample
    features features feature classification selection recognition accuracy use learned
    figure figure shown shows left right different system first
    function function functions value approximation given optimization cost defined
    gaussian gaussian covariance mean process posterior likelihood prior density
    gradient gradient stochastic optimization descent convergence methods convex gradients
    graph graph graphs edge edges nodes vertices random node
    image image images visual pixels pixel segmentation patches vision
    information information mutual entropy stimulus based analysis different processing
    input input output inputs units unit system patterns pattern
    kernel kernel kernels svm feature test support based hilbert
    latent latent variables variable variational observed inference tensor posterior
    layer layer layers deep hidden units first networks convolutional
    learning learning learn active machine rule learned examples weight
    Table 6: The top 20 topics, as fit by BRETT on the NeurIPS data set. The top 8 words (ranked by the entries in $\bm{\Phi}$) in each topic are shown, and topics are labelled by their associated anchor word. This is very often also the word with the highest weight in its respective topic. Topics are ranked according to the $\lambda$-criterion.
    topic 1 2 3 4 5 6 7 8
    topic 1 missing data erent values classi mask imputation cation
    topic 2 object objects shape figure contour surface recognition depth
    topic 3 direction head position activity cells location place spatial
    topic 4 protein alignment sequence sequences proteins species structure set
    topic 5 prediction data predictions predictive predict individual population predicted
    topic 6 log bound divergence lower exponential family exp upper
    topic 7 layer layers convolutional deep training network learning networks
    topic 8 estimate estimation estimator variance estimates bias estimated mean
    topic 9 learning active algorithm hypothesis examples query queries learner
    topic 10 video motion frame pose frames camera using human
    topic 11 label labels instances instance labeling labeled set positive
    topic 12 algorithm algorithms time method number problem iteration step
    topic 13 decision agents belief agent market price value utility
    topic 14 words word topic document documents topics text lda
    topic 15 training deep neural learning networks bengio layer trained
    topic 16 regret algorithm bandit arm problem arms bound time
    topic 17 ranking rank permutation pairwise top order set ranked
    topic 18 regression regularization lasso selection sparse regularized group norm
    topic 19 human figure participants subjects experiment people cognitive humans
    topic 20 gaussian covariance process function mean regression noise data
    Table 7: The “top” 20 topics, as fit by an STM on the NeurIPS data set. The top 8 words (ranked by the probability of each word within its respective topic) in each topic are shown. As no anchor words exist, topics are simply indexed by integers. Topics are not ranked in any sense; shown is simply the indexing discovered by the algorithm.

### v2

@@
 The issue of estimating covariate effects on topics in documents is motivated by questions of regional differences in the discourse of flavours of Canadian craft beers.
-grains are predominantly grown in the prairies and the freshness of the ingredients may result in beers that have grainier, maltier flavours.
-While most breweries will produce a suite of beer styles including India Pale Ales, Stouts, Pilsners, etc $\ldots$, we hypothesize that the flavours within those beer categories will tend towards regional effects that can be quantified from the text used to describe the beer flavours.
-Previous work on single malt scotch whiskies used presence or absence of specific terms as covariates (Lapointe and Legendre, 1994), however that dataset was produced by a single reviewer based on a standardized language.
-Our dataset is the unstructured text from thousands of different product reviewers, making binary covariates a less useful measure compared to covariates based on the strength of topics.
-The difficulties in including covariate effects manifest in two, related ways.
+Grains are predominantly grown in the prairies while hops are predominantly grown in Southern Ontario and South-Western British Columbia and interest lies in whether or not the proximity of ingredient production is associated with increased discourse in their flavours.
+Previous work on single malt scotch whiskies used presence or absence of specific terms as covariates to classify location of origin (Lapointe and Legendre, 1994), however that dataset was produced by a single reviewer based on a standardized language.
+Our interest is not in predicting the unknown location of origin of a product based on a description of its flavour, but rather we solve the tourist guidebook problem.
+Given a region, what should we expect of the flavour discourse about an unknown beer as measured in a particular attribute (topic) of interest.
+The difficulties in including covariate effects on topics of discourse manifest in two, related ways.
 First is the challenge of implementing bespoke approximate inference techniques, like Markov chain Monte Carlo (MCMC) and variational inference, that are needed to fit PTMs incorporating the desired extra structure.
-Fortunately, research in automating this step of the modelling pipeline has proliferated in the last decade; see, for example, advances in scalable variational inference (Kucukelbir et al., 2017; Ranganath et al., 2014) and their implementation in increasingly-popular probabilistic programming languages such as Stan (Carpenter et al., 2017).
+Fortunately, research in automating this step of the modelling pipeline has proliferated in the last decade; see, for example, advances in scalable variational inference (Kucukelbir et al., 2017; Ranganath et al., 2014) and their implementation in probabilistic programming languages such as Stan (Carpenter et al., 2017).
 These general systems allow users to specify broad classes of probabilistic models and treat inference as a black-box procedure whose details are hidden from the analyst.
@@
 With this flexibility comes computational efficiency, as efforts can be focused on statistics for which fast, reliable algorithms exist.
-Second, the variability of these statistics can be estimated using non-parametric resampling techniques like the bootstrap.
+Second, when the variability of these statistics may be estimated by likelihood means it can be estimated using non-parametric resampling techniques like the bootstrap.
 This philosophy of inference is illustrated by an analogy to standard OLS regression: one need not assume that the true data-generating mechanism is given by a Gaussian linear model for the OLS estimates to usefully summarize the data.
@@
 The paper is organized as follows.
-Section (2) establishes notation and lays out the general mathematical setup before providing a brief account of non-negative matrix factorization and its ability to find latent structure in text data.
+Section 2 establishes notation and lays out the general mathematical setup before providing a brief account of non-negative matrix factorization and its ability to find latent structure in text data.
 The section continues by outlining how separable NMF leads to the concept of anchor words and tractable algorithms that solve the NMF problem uniquely and globally.
-Section (3) leverages these developments in connecting NMF to ordinary regression models, permitting the inclusion of covariate effects (the motivating problem of this work).
-Statistical inference is addressed within this framework by showing how a bootstrap scheme can be added in a way that obviates repeatedly solving the NMF problem.
-Section (4) serves two purposes.
-First, BRETT is compared to the STM framework using publicly-available text data from NeurIPS conference papers.
-The section then applies BRETT to a data set of Canadian beer reviews, paying particular attention to regional differences as encoded by a covariate associated to each document.
-Conclusions and future research directions follow in section (6).
-2 Methods
+Section 3 leverages these developments in connecting NMF to ordinary regression models, permitting the inclusion of covariate effects (the motivating problem of this work).
+Statistical inference is addressed within this framework by showing how a beta regression or least squares with a bootstrap scheme can be added in a way that obviates repeatedly solving the NMF problem.
+Section 4 serves two purposes.
+First, BRETT is compared to the STM framework in inference and compute time using publicly-available text data from NeurIPS conference papers.
+The section then considers a simulation study showcasing how inferential accuracy evolves with the number of words per document and the number of documents.
+Section 5 applies BRETT to a data set of Canadian beer reviews.
+Hypotheses related to differences in discourse associated with beer styles are tested, followed by testing regional differences in flavour hypothesized to be associated with proximity to production of beer ingredients.
+This section ends with discussion about stability of BRETT with respect to the number of topics, the curation of the text corpus, and sampling variability of NMF.
+Conclusions and future research directions follow in section 6.
+2 Topic Modelling
 The term-document matrix (TDM) $\mathbf{X}\in\mathbb{N}^{V\times D}\subset\mathbb{R}_{+}^{V\times D}$ contains entries $x_{ij}$ representing the counts of word $i$ appearing in document $j$ for a corpus of $V$ unique words and $D$ documents.
 Note that substantial pre-processing is often required to transform a raw corpus into a form amenable to analysis.
-As an auxiliary piece of data, let $\mathbf{Z}\in\mathbb{R}^{D\times P}$ be a model matrix encoding document-level covariates (these can be discrete or continuous) and an intercept term.
+As an auxiliary piece of data, let $\mathbf{Z}\in\mathbb{R}^{D\times P}$ be a model matrix encoding document-level covariates and an intercept term.
 The only stochastic assumption we employ is that
 $(\mathbf{X},\mathbf{Z})\sim\pi(\mathbf{X},\mathbf{Z})$
-where $\pi$ is some fixed but unknown probability measure.
-Formally, this measure is required to be sufficiently well-behaved so as to permit unproblematic bootstrapping.
+where $\pi$ is some fixed but unknown probability measure. Formally, this measure is required to be sufficiently well-behaved so as to permit unproblematic bootstrapping.
 For standard covariate-free topic modelling, one could compute a statistic
@@
 $\psi_{2}\left(\mathbf{X},\mathbf{Z}\right),$
-serves the same purpose but also incorporates $\mathbf{Z}$ in an appropriate way.
-Finally, estimate the sampling distribution of these statistics by bootstrapping or other means.
-This sec constructs specific instances of $\psi_{1}$ and $\psi_{2}$, as well as efficient methods of estimating their sampling distributions.
+serves the same purpose but also incorporates $\mathbf{Z}$ in an appropriate way. Finally, estimate the sampling distribution of these statistics by bootstrapping or other means.
 2.1 Non-negative Matrix Factorization
-The essential feature of NMF is its attempt to factor $\mathbf{X}$ into two “simpler” matrices $\bm{\Phi}$ and $\bm{\Theta}$, also containing non-negative entries,
+NMF factorizes $\mathbf{X}$ into two “simpler” matrices $\bm{\Phi}$ and $\bm{\Theta}$, also containing non-negative entries,
 $\mathbf{X}\approx\bm{\Phi}\bm{\Theta}$
@@
 Treating (2.1) as defining statistics requires appealing to the concept of anchor words, central in formulating separable NMF.
-First studied by Donoho and Stodden (Donoho and Stodden, 2004), the separability assumption states that within ${\bm{\Phi}}$ lies a $T\times T$ diagonal matrix, possibly after permuting ${\bm{\Phi}}$’s rows appropriately.
+The separability assumption states that within $\bm{\Phi}$} lies a $T\times T$ diagonal matrix, possibly after permuting $\bm{\Phi}$}’s rows appropriately (Donoho and Stodden, 2004).
+Separability means that for each column $j$ we can find an entry $\phi_{ij}>0$ such that $\phi_{kj}=0$ for all $k\neq i$.
+Such a word is called an anchor word, for it “ties down” a particular topic.
 In the language of topics, this means that for each topic (column of $\bm{\Phi}$}) there exists a word with non-zero weight only within that topic.
-More formally, for each column $j$ we can find an entry $\phi_{ij}>0$ such that $\phi_{kj}=0$ for all $k\neq i$.
-Such a word is called an anchor word, for it “ties down” a particular topic.
+For separability to be valid, $X$ must be non-negative and there must be data points spread across $T$ faces of the positive orthant, so that $X$ has at least a rank $T$ orthogonal decomposition into $T$ topics.
+Any corpus of documents differing in discourse and meaning, implying a sparse TDM, will meet the separability assumption for values of $T$ which are much larger than necessary for inference or data summarization.
 Anchor words have been studied and used extensively in developing provably efficient algorithms for estimating PTMs (Arora et al., 2012; Arora et al., 2018; Gillis, 2014).
@@
 $\text{subject to}\quad\hat{\mathbf{Y}}\succeq 0.$ (3)
-Solving (2.2) amounts to a series of non-negative least squares (NNLS) problems, one for each of the $V-T$ non-anchor rows in $\mathbf{X}$.
+Solving (2.2) amounts to a series of non-negative least squares (NNLS) problems, one for each of the VTV-T non-anchor rows in $\mathbf{X}$.
 NNLS is well-studied and – crucially – convex (Chen and Plemmons, 2009); this establishes the overall convexity of (2.1) with the additional anchor word constraint.
@@
 2.3 Determining Anchor Words
-The above discussion assumes anchor words are known in advance. In practice, they must be found or – as is done here – choosen.
+The above discussion assumes anchor words are known in advance. In practice, they must be found or – as is done here – chosen.
 In some circumstances the analyst may have strong interests in using hand select anchors for subsequent inference.
-In our experience, a better approach is to employ a data driven tool for procuring anchors.
+In our experience, a better approach is to employ a data driven tool for procuring anchors which are then assessed for their interpretation in measuring hypothesized relationships.
 Topics represent a set of potential features from which the analyst can choose as interpretable, inferential measures.
@@
 While this may seem unmotivated, note that it is actually crucial to interpretability.
-Elaboration of this point is deferred to section (3).
+Elaboration of this point is deferred to section 3.
 SPA can be informally described by considering the $V$ rows of $\mathbf{X}$ as vectors in $\mathbb{R}^{D}$.
-  • Take the first anchor to be the point farthest from the origin; by the definition of the 2-norm, this is just the word that is most often within documents in the corpus.
+  • Take the first anchor to be the point farthest from the origin; by the definition of the 2-norm, this is the word that is most often within documents .
 Note that non-informative “stopwords” are removed as a pre-processing step prior to constructing $\mathbf{X}$.
@@
 This could be a genuine problem in corpora with very small documents, but experience shows that in reasonably sized collections (such as the ones we study later) these “empty” documents are rare.
-Alternatively, the raw vocabulary could be transformed into BERT word embeddings Devlin et al.
-(2019) before constructing the TDM Dieng et al.
-(2020).
+Alternatively, the raw vocabulary could be transformed into BERT word embeddings (Devlin et al., 2019) before constructing the TDM (Dieng et al., 2020).
+We also note that anchor word selection is subject to sampling variability. This is discussed in Section 5.6 with particular attention to the real data application.
 Input: a TDM $\mathbf{X}$.
+2mm
   • Compute the QR decomposition of $\mathbf{X}^{\mathrm{T}}$ (with column pivoting).
   • Extract the anchor indices $\{a_{1},a_{2},\ldots,a_{T}\}\subset\{1,2,\ldots,V\}$ from the pivoting strategy.
-[…]
+  • Form $\mathbf{X}_{\dagger}$ from the rows of $\mathbf{X}$ coinciding with the anchor indices.
+Denote the remaining matrix of non-anchor rows as $\mathbf{X}_{\ddagger}$.
+Find $\hat{\mathbf{Y}}^{\text{opt.}}=\arg\underset{\mathbf{Y}\succeq 0}{\min}\hskip 5.69054pt{\left\lVert\mathbf{X}_{\ddagger}-\mathbf{Y}\mathbf{X}_{\dagger}\right\rVert}_{\mathrm{F}}$ by solving VTV-T NNLS problems.
   • For $j\in\{1,2,\ldots,T\}$, compute $\hat{\lambda}_{j}=1\big/\left(1+\sum_{i=T+1}^{V}\hat{y}_{ij}^{\text{opt.}}\right)$.
@@
     $\bm{\Phi}$, by concatenating $\bm{\Lambda}$ and $\bm{\Gamma}$ column-wise.
-    Output: $\psi_{1}=\left\{\hat{\bm{\Phi}},\hat{\bm{\Theta}}\right\}$.
+  • 2mm Output: $\psi_{1}=\left\{\hat{\bm{\Phi}},\hat{\bm{\Theta}}\right\}$.
+    2mm
     Algorithm 1 NMF-based topic modelling with anchor words
@@
     In unsupervised methods like clustering and topic modelling, there is no ground truth with which to compare.
-Several measures have been proposed to assess the number and quality of topics.
-Perplexity measures how well a held out set of documents conform to the topic model, however it has been shown to correlate poorly with human interpretable topics (Chang et al., 2009).
-UMass coherence considers the frequency of occurence of within-topic word pairs as a measure of the unique interpretability of the topic construction as assessed by the log ratio of the number documents in which the most important words co-occur compared to number of documents containing a word (Mimno et al., 2011).
-Variations attempt to improve the measure’s ability to capture what humans consider to be good topics (Newman et al., 2010; Röder et al., 2015).
+Several measures have been proposed to assess the number and quality of topics including Perplexity(Chang et al., 2009), UMass coherence (Mimno et al., 2011) and variations attempting to improve the measure’s ability to capture what humans consider to be good topics (Newman et al., 2010; Röder et al., 2015).
 However, it is notable that the target of these methods remains the ability for a human to interpret topics and consider them to be appropriate as compared to domain expertise.
@@
 As with any quantitative discipline, the scientist must determine if the metric is appropriate and targets, at least in proxy, the analytic goal.
-    Well curated data that is specific to the analytic goal should be combined with a moderate number of topics relative to the number of documents. This is explored empirically in Section 5.4.
+Well curated data that is specific to the analytic goal should be combined with a moderate number of topics relative to the number of documents.
+Using the context of the real data application, a comparison of using a general or curated corpus is discussed in Section 5.5, the impact of the number of topics is discussed in Section 5.4, and stability with respect to anchor selection is discussed in Section 5.6.
     3 Modelling Associations Between Topics and Covariates
@@
 Modelling focuses on assessing the extent to which covariates affect $\bm{\Theta}$.
-This section outlines a fast OLS routine for the special case of categorical covariates and then continues into more realistic regression models for continuous covariates.
+This section outlines a fast OLS routine for the special case of categorical covariates and then continues into more general Beta regression models.
     When cast as a PTM, categorical covariates in an Ordinary Least Squares (OLS) hold the interpretation that coefficients are the estimated deviation of the mean in moving in or out of a category.
 Maintaining this interpretability through confidence intervals necessitates the use of bootstrap confidence intervals outlined in Section 3.1.
-This is expanded into the general purpose Beta-regression approach outlined in Section 3.2.
+This is expanded into the general purpose Beta-regression approach in Section 3.2.
     3.1 OLS Regression and Bootstrapping
-    Focusing on $\bm{\Theta}$ (as extracted by algorithm (1)) with elements $\theta_{ij}$ describing the probability of topic $i$ in document $j$ and covariate vector $\mathbf{z}^{j}$ is the $j^{\text{th}}$ row of $\mathbf{Z}$, write
+    Focusing on $\bm{\Theta}$ as extracted by algorithm 1, with elements $\theta_{ij}$ describing the probability of topic $i$ in document $j$ and covariate vector $\mathbf{z}^{j}$, the $j^{\text{th}}$ row of $\mathbf{Z}$, write
     $\widetilde{\theta}_{ij}\approx\langle\mathbf{z}^{j},\bm{\beta}_{i}\rangle,$
@@
     $\text{minimize}\quad{\left\lVert\widetilde{\bm{\Theta}}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$ (4)
-    The constraint $10<\theta_{ij}<$ 1 imposes heteroskedasticity and directionality to the residuals of (4).
-The breach of the Gauss-Markov assumptions degrades the quality of OLS interval estimates.
-Instead bootstrap intervals are used.
-    The decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$ where $\left(\bm{\Lambda}\right)^{-1}$ is diagonal scales the rows of $\mathbf{X}_{\dagger}$, an effect that is undone by row normalization.
+    The constraint $0<\theta_{ij}<1$ imposes heteroskedasticity and directionality to the residuals of (4). The breach of the Gauss-Markov assumptions necessitates the use of bootstrap intervals.
+    The decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$, where $\left(\bm{\Lambda}\right)^{-1}$ is diagonal, scales the rows of $\mathbf{X}_{\dagger}$, an effect that is undone by row normalization
     Thus, (4) can be reformulated as
-    $\text{minimize}\quad{\left\lVert\widetilde{\mathbf{X}}_{\dagger}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$
+    $\text{minimize}\quad{\left\lVert\widetilde{\mathbf{X}}_{\dagger}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$ (5)
     Crucially, this only involves the scaled TDM.
-The upshot is that algorithm (1) need not be re-run when using resampling methods to estimate the sampling distribution of $\mathbf{B}$ – nowhere does it depend on $\bm{\Theta}$.
-    Algorithm (2) describes a bootstrapping procedure that takes advantage of this simplification.
+The upshot is that algorithm 1 need not be re-run when using resampling methods to estimate the sampling distribution of $\mathbf{B}$ – nowhere does it depend on $\bm{\Theta}$.
+Algorithm 2 describes a bootstrapping procedure that takes advantage of this simplification.
 Resampling the documents, columns of $\mathbf{X}$, allows efficient estimation of the sampling distribution of $\psi_{2}$.
     Input: Observed TDM $\mathbf{X}$, number of bootstrap samples $b$, anchor indices $\{a_{1},a_{2},\ldots,a_{T}\}$ (as found by algorithm (1) or some other method).
+    2mm
     • For $i\in\{1,2,\ldots,b\}$:
@@
         Find $\mathbf{B}^{i}=\arg\underset{\mathbf{B}}{\min}\hskip 5.69054pt{\left\lVert\widetilde{\mathbf{X}}_{\dagger}^{S}-\mathbf{B}^{\text{T}}\left(\mathbf{Z}^{S}\right)^{\text{T}}\right\rVert}_{\mathrm{F}}$ by solving $T$ OLS problems.
-    Output: $\{\psi_{2}^{i}=\mathbf{B}^{i}\}_{i=1}^{B}$.
+    2mm Output: $\{\psi_{2}^{i}=\mathbf{B}^{i}\}_{i=1}^{B}$.
+    2mm
     Algorithm 2 BRETT: Bootstrapped/Beta Regression Effects Topic Trends.
@@
 While we find that using SPA to pre-select anchor words works well in practice, it could be altered with a bespoke tool without breaking the essential inferential mechanism.
-So long as “good” anchor words are produced (in the sense that the analyst finds them an informative summary of their corpus), useful for hypothesis testing and effect estimation can be devised without re-computing anchor words at each bootstrap iterate.
+So long as “good” anchor words are produced (in the sense that the analyst finds them an informative summary of their corpus), useful hypothesis testing and effect estimation can be devised without re-computing anchor words at each bootstrap iterate.
     3.2 Beta Regression
-    Again focusing on attention on $\bm{\Theta}$ (as extracted by algorithm (1)), recall that the entries $\theta_{ij}\in[0,1]$ describes the probability (or “weight”) assigned to topic $i$ in document $j$.
-After selecting topic $i$ of interest to the analyst, and normalizing $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{i}\theta_{ij}$ so that rows sum to 1, vector $\widetilde{\theta}_{i}$ contains elements interpretable as the probability a word from document $j$ as being from topic $i$.
-For a fixed $i$, modelling $\theta_{ij}$ as depending on a vector of covariates $\mathbf{z}^{j}$ through coefficient vector $\bm{\beta}_{i}\in\mathbb{R}^{P}$,
-    $\widetilde{\theta}_{ij}\sim Beta(\mbox{mean}=\mu,\mbox{precision}=\sigma),with$
-    $g_{\mu}(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$
-    The mean-precision parameterization of the Beta provides a more intuitive interpretation of regression (Vasconcellos and Cribari-Neto, 2005).
-The link functions, $g_{\mu}(\mu)$ and $g_{\sigma}(\sigma)$, provide mappings from potentially different subsets of $\mathbf{z}^{j}$ into the appropriate spaces with covariate driven variation (Simas et al., 2010).
+    Again focusing on attention on $\bm{\Theta}$, as extracted by algorithm 1, recall that the entries $\theta_{ij}\in[0,1]$ describe the probability (or “weight”) assigned to topic $i$ in document $j$.
+After selecting topic $i$ of interest to the analyst, and normalizing $\widetilde{\theta}_{ij}=\theta_{ij}/\sum_{i}\theta_{ij}$ so that rows sum to 1, vector $\widetilde{\theta}_{i}$ contains elements interpretable as the probability of a word from document $j$ as being from topic $i$.
+Again, the decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$, followed by the rescaling in $\widetilde{\theta}$ highlights that Beta regression is conditional only on the anchor block.
+For a fixed $i$, modelling $\widetilde{\theta}_{ij}$ as depending on a vector of covariates $\mathbf{z}^{j}$ through coefficient vector $\bm{\beta}_{i}\in\mathbb{R}^{P}$,
+    $\widetilde{\theta}_{ij}\sim Beta(\mbox{mean}=\mu,\mbox{precision}=\sigma).$ (6)
+    The mean-precision parameterization of the Beta regression model provides a more intuitive interpretation of regression (Vasconcellos and Cribari-Neto, 2005).
+The link functions map potentially different subsets of $\mathbf{z}^{j}$ into the appropriate spaces with covariate driven variation (Simas et al., 2010)
+    $g_{\mu}(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$ (7)
     Beta regression can be applied directly to this problem conditional on the extraction of $\bm{\Theta}$.
@@
 The first analysis considers the time evolution of discourse in papers from the NeurIPS conference to compare STM and BRETT.
-BRETT is shown to be much faster than the STM framework while retaining inferential accuracy.
-The second analysis considers the regional differences in vocabulary used to describe beer flavours.
+A simulation example follows showcasing behaviour with respect to changes in document length and number.
     4.1 Comparisons with STMs
@@
 One could compute regression-style statistics using $\bm{\Phi}$ (akin to what was accomplished in the previous section) to mimic this ability, but the efficient bootstrap-based error bars would be lost, as one would need to re-compute $\bm{\Phi}$ with each bootstrap sample.
-Accordingly, we only compare to those aspects of STMs most directly comparable to BRETT: their topic discovery and incorporation of “$\bm{\Theta}$-influencing” covariates.
+Accordingly, we only compare to those aspects of STMs most directly comparable to BRETT: their topic discovery and incorporation of $\bm{\Theta}$-influencing” covariates.
     Papers from the NeurIPS conference between the years 1987 and 2015 are used as a toy example.
@@
 First, anchor words provide “labels” of each topic, though this point seems moot in this example: the topics found by both methods appear easily understandable to humans.
-More importantly, BRETT ranks topics using the $\lambda$-criterion described in section (LABEL:sec:nmf), a feature lacking in PTMs that emphasizes how predictive the anchor is of its constituent words.
+More importantly, BRETT ranks topics using the $\lambda$-criterion described in section 2.2, a feature lacking in PTMs that emphasizes how predictive the anchor is of its constituent words.
 BRETT ranks specific, semantically interesting topics higher.
-    Although the topics differ outside of their leading words, this ensures that they are at least somewhat semantically comparable; in the case of the “Gaussian” topic, tables (6) and (7) in appendix A suggest that both methods are referring primarily to Gaussian processes.
-    Data fits are shown in Figure 1.
+    Although the topics differ outside of their leading words, this ensures that they are at least somewhat semantically comparable; in the case of the “Gaussian” topic, tables 7 and 8 in appendix A suggest that both methods are referring primarily to Gaussian processes.
+    Regression fits to the data for the are shown in Figure 1 using binned time as the discrete regression covariates.
 The STM results are from OLS and Beta regression as estimated by the mean of 10,000 draws from the variational posterior and computing the associated regression coefficients each time.
-The BRETT results are from fitting Beta regression directly and from the mean of 10,000 bootstrap OLS samples.
+The BRETT results are shown fitting Beta regression directly and also from the mean of 10,000 bootstrap OLS samples.
 Results are similar owing to the discrete nature of the covariates.
-    Uncertaintly characterization differs in interpretation across the methods.
-STM tools and BRETT OLS sample the regression surface producing uncertainty intervals for the model fit.
-BRETT Beta instead fits a Beta distribution to the data, naturally producing prediction intervals for new observations.
-This difference characterizes the differing widths of intervals overlayed in Figure 1.
+    Uncertainty characterization differs in interpretation across the methods.
+STM and BRETT OLS sample the regression surface producing uncertainty intervals for the model fit.
+BRETT Beta regression instead fits a Beta distribution to the data, producing prediction intervals for new observations.
+This difference in interpretation is highlighted in the differing widths of intervals overlayed in Figure 1.
     (a) topic ”gaussian”
@@
 The number of topics $T$ was chosen to range over the values $\{10,50,100,200\}$ and a binary variable $C\in\{T,F\}$ was devised, representing whether covariates should be included in the analysis or not.
-If not, classical covariate-free topic modelling was performed.
+If not, classical covariate-free topic modelling was performed where STM and BRETT reduce to LDA with NMF respectively.
 Otherwise, covariate effects were estimated and 1000 simulations computed to assess the uncertainty in these estimates.
@@
 All $4\times 2=10$ unique settings of $T$ and $C$ were timed using the same software on the same machine; figure (2) shows the results.
-BRETT is typically an order of magnitude faster and incurs minimal slowdowns when incorporating the extra regression functionality.
+BRETT and NMF are typically an order of magnitude faster and incurs minimal slowdowns when incorporating the extra regression functionality.
     Figure 2: Running times for BRETT and STMs with covariates (left) and without (right).
-    4.1.3 Inferential Accuracy
+    4.2 Inferential Accuracy
     Beta regression is known to perform well for estimating covariate effects (Simas et al., 2010), consequently this simulation study focuses on the sampling variability of the NMF decomposition and it’s propagation through the regression procedure.
@@
 Given the simulation mechanism, this study explores the variability with respect to changes in $N_{D}$, and $D$ with fixed anchors, using both recalculated $\Phi$ specific to each simulation or fixing $\Phi$ at the value used to simulate the TMDs.
+Recalculating $\Phi$ for each dataset is equivalent to having an analyst select the anchor words and let the data fine tune the interpretation of the topic.
+Holding $\Phi$ fixed at the value used to construct the simulation defines $\Theta=(\Phi^{\prime}\Phi)^{-1}\Phi^{\prime}X$ and is equivalent to the analyst pre-defining topics from an external independent dataset and uses those topic definitions to define document topic weights.
     To ensure a consistent simulation setup, the simulation details are somewhat atypical.
@@
 To assess the stability of the regression coefficient estimates, a pseudo-ground truth must be devised.
-A pseudo ground truth is constructed by summing the 1000 TDMs elementwise into the equivalent of a TMD where each document has $1000\times N_{D}$ words per document.
-    The alternatives were to elimiate low probability words resulting in a much smaller vocabulary or allow a varying vocabulary size across simulations.
+A pseudo ground truth is constructed by summing the 1000 TDMs element-wise into the equivalent of a TMD where each document has $1000\times N_{D}$ words per document.
+    The alternatives were to eliminate low probability words resulting in a much smaller vocabulary or allow a varying vocabulary size across simulations.
 These alternative simulation designs are also expected to alter the covariate effects from their simulated targets.
 The Mean Squared Error for the covariate effect effect is shown in figure 3.
+The strategy of recalculating $\Phi$ for each TDM while holding the anchors fixed performs well with the MSE decreasing quickly in both number of documents and number of words per document.
+Allowing $\Phi$ to be recalculated introduces some variability in the balance of words within topic, but overall seems to estimate a $\Phi$ that is close to its asymptotic baseline value.
+The strategy of recalculating $\Phi$ for each dataset results in around an order of magnitude decrease in MSE compared to using a fixed $\Phi$.
+In the fixed $\Phi$ strategy, $\Phi$ is held at the wrong value as it does not account for the simulation mechanism that ensures the vocabulary size remains fixed leading to a decrease in performance.
+This suggests that the analyst is best served by selecting anchor terms and letting the NMF routine define the allocation of terms into topics.
     Figure 3: Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample (left) or using a single fixed $\Phi$ for all simulations (right).
-Mean Squared Error in estimating regression effects using Beta regression with respect to the number of documents (rows) and number of words per document while recalculating $\Phi$ for each sample.
     5 Application: The “Beer Data”
-    A dataset about Canadian beer reviews is detailed in section 5.1 is used to show the method is able to extract well known results about flavours that define different beer styles in section 5.2.
+    Collection and processing of the Canadian beer review dataset is detailed in section 5.1 and used to show BRETT is able to extract well known results about flavours that define different beer styles in section 5.2.
 Section 5.3 tests hypotheses about whether local production of key ingredients has an affect on discourse around beer flavours.
@@
     5.2 Beer Style Hypotheses
-    The dataset was filtered down to 7 beer styles, counts of which are given in table 1. The NMF model was run with 50 topics.
+    The dataset was filtered down to 7 beer styles, counts of which are given in table 1. NMF was run with 50 topics.
     The ability to test hypotheses depends on having a good metric for the characteristic of interest.
-The 50 topics, were examined manually to find measures of flavour attributes that coincide with the characteristics of the beer style definitions (Palmer, 2017).
+The topics were examined manually to find measures of flavour attributes that coincide with the characteristics of the beer style definitions (Palmer, 2017).
 Some flavour attributes were spread into multiple topics, similar to the way lab sciences often indirectly measure attributes of interest from different perspectives and experiments.
@@
     • Saisons are have citrus and peppery tones with considerable flavour influenced from the yeast (H8).
-    The main terms composing the topics to test hypotheses are given in table 2. For the sake of brievity the table shows the most relevant terms from the top 20 most important terms per topic.
-    The hypotheses are tested through logistic link function in Beta regression.
+    The main terms composing the topics to test hypotheses are given in table 2.
+For the sake of brevity the table shows the most relevant terms per topic.
+The hypotheses are tested using Beta regression from equation (6) where the design matrix $Z\in\{0,1\}^{D\times 7}$ is defined with categorical covariates corresponding to the 7 beer styles through link functions
+    $g_{\mu}(\mu)=logit(\mu)=\mathbf{B}_{\mu}^{\text{T}}\mathbf{z}^{j},\mbox{ and }g_{\sigma}(\sigma)=logit(\sigma)=\mathbf{B}_{\sigma}^{\text{T}}\mathbf{z}^{j}.$ (8)
+    In each case one beer style is chosen as a baseline (intercept) and all other covariates are interpreted as deviations thereof for testing the null hypothesis of no difference in discourse about a flavour characteristic compared to the baseline beer style.
     Regression coefficients for hypotheses (H1-H8) are shown in table 3 as the deviation from the baseline beer style.
 In all cases our choice of baseline has a strong topic effect.
-Deviations from the baseline are substantial and negative except in ways that coincide with expected definitions and closely related beer styles.
+Deviations from the baseline are substantial and negative except in ways that coincide with expected definitions and closely related beer styles suggesting a reduction in discourse associated with changing beer style.
 Imperial IPAs are considerably stronger than IPAs and often darker and with more hops.
@@
 The dark malt base provides a creaminess, often enhanced by oatmeal (in stouts in particular), or the addition of nitrogen carbonation.
-Pumpkin ales also have creaminess from their pie-like flavours appearing as similarity to baselines (H7 difference between stouts and porters p-value .26, and difference between stouts and pumpkin ales p-value 0.070).
+Pumpkin ales also have creaminess from their pie-like flavours resulting in a similarity with baselines (H7 difference between stouts and porters p-value .26, and difference between stouts and pumpkin ales p-value 0.070).
 Saisons are considerably yeastier in flavour than other beers, the closest being lagers and pumpking ales (H8 difference between saisons and lagers p-value 2e-36, and difference between saisons and pumpkin ales p-value 8e-10).
 Lagers are generally light in flavour and sometimes the yeast is a dominant flavour, but far from the prominance of a saison.
-Pumpkin ales have considerable variability spanning and could be modified from anything from a saison to a stout base.
-These effects are all reflected in the prediction intervals of figure 4 for topics in table 2.
-Beta Distributions have full support over $[0,1]$ leaving effect size prediction intervals wide with some strong prediction confidence close to 0, even when an overall effect is large as is seen in all the baseline beer styles for the hypotheses.
+Pumpkin ales have considerable variability as they could be modified from nearly any beer style.
+These effects are all reflected in the fitted distribution of figure 4 for topics in table 2.
     topic defining terms
-    grapefruit pineapple, fresh, crisp, love, citrusy, backbone, excellent, delicious, juicy, lacing, resiny, clean
-    tropical mango, pineapple, fruits, lemon, gold, tangerine, dank, juicy, rind, citra, peach, papaya
-    pine piney, resin, fresh, hoppy, balance, oily, west coast, balanced, huge, sticky, plenty, dipa
-    yeast spice, white, pepper, lemon, spicy, wheat, golden, belgian, yellow, yeasty, banana, fruity
-    brett funk, tart, funky, oak, sour, wine, tartness, lemon, apple sourness, barnyard, wild
-    pumpkin cinnamon, spices, nutmeg, spice, pie, ginger, clove, amber, allspice, copper, fall, flavour
-    roasted black, burnt, roasty, oatmeal, tan, thick, smooth, oats, sweetness, opaque, brown, espresso
-    coffee espresso, strong, beans, brown, mouthfeel, opaque, drinkability, bodied, heavy, hints, retention, barley
-    chocolate cocoa, milk, mocha, hint, black, smooth, mouthfeel, roast, characters, delicious, lacing, chocolatey
-    creamy smooth, thick, brew, excellent, rich, cream, bodied, dense, sweetness, nitro, silky, feel
+    tropical mango, pineapple, fruits, lemon, gold, tangerine, dank, juicy, rind, citra, peach
+    grapefruit pineapple, fresh, crisp, love, citrusy, backbone, excellent, delicious, juicy, lacing, resiny
+    pine piney, resin, fresh, hoppy, balance, oily, balanced, huge, sticky, plenty
+    roasted black, burnt, roasty, oatmeal, thick, smooth, oats, sweetness, opaque, brown, espresso
+    coffee espresso, strong, beans, brown, mouthfeel, opaque, drinkability, bodied, heavy, hints, barley
+    chocolate cocoa, milk, mocha, hint, black, smooth, mouthfeel, roast, characters, lacing, chocolatey
+    creamy smooth, thick, brew, rich, cream, bodied, dense, sweetness, nitro, silky, feel
+    yeast spice, white, pepper, lemon, spicy, wheat, golden, belgian, yellow, yeasty, banana
     Table 2: Topics used to test flavour attribute hypotheses for different beer styles. The defining words were chosen from the top 20 most predictive terms within topics.
     Hypothesis Imperial IPA IPA Lager Porter Pumpkin Ale Saison Stout
-    H1 - tropical 0.28 NS -2.77$\star$ -3.07 -4.19 -4.90 -1.39 -3.75
-    H2 - grapefruit -0.15 NS -2.75$\star$ -2.57 -4.94 -4.27 -2.20 -8.39
-    H3 - pine 0.01 NS -3.35$\star$ -2.56 -4.52 -5.48 -2.06 -4.41
-    H4 - roasted -5.10 -4.10 -3.95 -0.18NS -2.36 -3.46 -2.08$\star$
-    H5 - coffee -6.17 -4.53 -5.05 -0.42 -3.83 -4.93 -2.48$\star$
-    H6 - chocolate -6.32 -4.84 -6.77 -0.02NS -3.03 -4.89 -2.83$\star$
-    H7 - creamy -1.05 -0.92 -1.59 -0.22NS -0.62NS -1.38 -3.36$\star$
-    H8 - yeast -2.21 -2.00 -1.95 -2.68 -2.26 -1.35$\star$ -2.88
+[…]
+    H2 - grapefruit -0.15 NS -$\star -2.57$ -4.94 -4.27 -2.20 -8.39
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
     Table 3: Beta regression coefficients for hypotheses 1-8.
@@
     5.3 Regional Differences in Flavour Characteristics of IPAs
-    Interest lies in assessing if there are effects of ingredient production within the flavour category.
-Specifically we are interested in knowing if the Prairies have grainier, breadier IPAs or if BC and Ontario have hoppier IPAs owing to localized production of base ingredients.
-Hypothesis tests for regional effects on flavour discourse will be confounded with differences in the types and numbers of beers produced.
-Focusing within a single product category, IPAs are used to test hypotheses that there are regional differences in flavour discourse.
-The NMF model is refitted using 776 IPAs.
-Using too general a corpus of all beer styles to devise topics results in messier topics that are less interpretable or useful (see for example Tshitoyan et al.
-(2019), Luca et al.
-(2022)).
-NMF was fitted with 50 topics resulting in topics that interpretably showcase the flavours of interest for these hypotheses.
-    The grainy, bready flavour set can be assessed using earthy (H9) and caramel (H10) topics while the hoppier characteristics were assessed using grapefruit (H11) and tropical (H12).
+    Interest lies in assessing if there are geographic effects relating to proximity of ingredient production within their flavour categories.
+Specifically we are interested in knowing if the Prairies have grainier, maltier IPAs or if BC and Ontario have hoppier IPAs owing to localized production of base ingredients.
+Hypothesis tests for regional effects on flavour discourse will be confounded with differences in the types and numbers of beers produced, so we focus within a single product category, IPAs are used to test hypotheses.
+NMF is refitted using 776 IPAs (and Imperial IPAs) with 50 topics.
+    The grainy and malty flavours can be assessed using earthy (H9) and caramel (H10) topics while the hoppier characteristics are assessed using grapefruit (H11), tropical (H12), and pine (H13).
 Table 4 provides characteristic terms within these topics.
-    The probability of topic $Y_{i}$ for the $i^{th}$ beer is modelled with an intercept effect for the region of interest with other covariates representing the $j^{th}$ regional deviation from the intercept.
-    Regression coefficients for hypotheses (H9-H12) are shown in table 4.
-The Prairies show significantly more earthy and caramel flavour discourse than any other region.
-These topics represent flavours that are not typical of the IPA style, suggesting an influence of locally produced ingredients.
-The grapefruit and tropical topics describe different dimensions of hop flavours characteristic of IPAs.
-The vast majority of beers are brewed with dried, vacuum sealed hops that travel well with reasonable shelf life.
-It should be expected that discourse around all IPAs mention these topics with considerable probability.
-It is not surprising that there are almost no significant regional differences from the BC baseline.
+    Hypotheses are tested through logistic link function in Beta regression using equations (6) and (8) where the design matrix $Z\in\{0,1\}^{D\times 5}$ is defined with categorical covariates corresponding to the 5 regions.
+In each case one region style is chosen as a baseline (intercept) and all other covariates are interpreted as deviations thereof associated with changing regions.
+Under the null hypothesis of no difference between regions, we expect to observe a significant effect for the baseline while coefficients for all other regions take values of approximately 0.
+    Regression coefficients for hypotheses (H9-H13) are shown in table 5.
+The Prairies show significantly more earthy and caramel topic discourse than any other region.
+These topics represent flavours that are not typical of the IPA style, potentially influenced by local production of ingredients.
+The grapefruit, tropical, and pine topics describe different dimensions of hop flavours characteristic of IPAs.
+Almost no significant regional differences from the BC baseline were found.
 The exception is that Ontario is showing a somewhat positive difference from BC (H11 difference between BC and Ontario p-value 0.019) in discourse of the grapefruit topic characterized by crisp, citrusy, resiny terms.
-Prediction intervals for effects are shown in figure 5.
-    5.4 Stability With Respect to Number of Topics
-    When the procedure in Section 5.2 was performed using 150 topics, the results were largely the same since additional anchors extend the topics into a higher dimensional orthogonal space.
-The utility of new topics starts to break down when too many topics are added because each new topic chips away at existing topics.
-Eventually a topic may split into two, diffusing the covariate of interest and rendering the hypothesis more challenging to define.
-Note that the probability of a topic for a document is monotone decreasing in the number of topics meaning that results are conditional on the model structure.
-    5.5 Stability With Respect to General vs Curated Corpus
-    When the topics in Section 5.3 were formed using the entire beer dataset, the topics were tuned towards separating out the broad differences, mostly driven by different beer styles.
-When using those topics and subsetting the documents to just IPAs, the flavour effects of interest were too diffuse to be usefully measured.
-Language models have been shows to have the best success when trained on data tuned to the task of interest (Tshitoyan et al., 2019).
+The vast majority of beers are brewed with dried, vacuum sealed hops that travel well with reasonable shelf life explaining the lack of overall flavour effects associated with proximity to hop farms.
+Predicted distribution of topic discourse for effects are shown in figure 5.
+    Figure 5: Predicted distributions for flavour topics by region.
     topic defining terms
@@
     tropical mango, dank, yellow, rind, tangerine, papaya, citra, juicy, resin, golden
-    Table 4: Topics used to test flavour attribute hypotheses H9-H12 by region within the IPA beer style.
+    pine piney, resin, balance, huge, hoppy, oily, rind, balance, complex, lingering
+    Table 4: Topics used to test flavour attribute hypotheses H9-H13 by region within the IPA beer style.
     Hypothesis BC Prairie Ontario Quebec Atlantic
@@
     H12 - tropical -3.07 $\star$ -0.19NS 0.08 NS -0.12 NS 0.33 NS
+    H13 - pine -3.84 $\star$ -0.14NS 0.03 NS -0.34 NS -0.13 NS
     Table 5: Beta regression coefficients for hypotheses H9-H12.
@@
 Coefficients that are not significantly different from 0 at the 5% level are labelled with NS.
-    Figure 5: Prediction interval for flavour topic probability for different regions using 50 topics.
-    6 Conclusion
+    5.4 Stability With Respect to Number of Topics
+    When the procedure in Section 5.2 was performed using 150 topics, the results were largely the same since additional anchors project the documents into a higher dimensional orthogonal space.
+The utility of new topics starts to break down when too many topics are added.
+Eventually a useful topic will split into two, diffusing the covariate of interest and rendering the hypothesis more challenging to define.
+Note that the probability of a topic for a document is monotone decreasing in the number of topics meaning that results are conditional on the model structure.
+    5.5 Stability With Respect to General vs Curated Corpus
+    When the topics in Section 5.3 were formed using the entire beer dataset rather than the IPA subset, the topics that emerged were tuned towards separating out the broad style differences since these captured most of the variation in the dataset.
+The resulting flavour topics were overly general when a small number of topics was used and when larger numbers of topics were used, topics of interest fractured off of the general topics in such a way that necessitated a larger number of hypotheses to test for the flavours of interest.
+Language models have similarly been shown to have the best success when trained on data tuned to the task of interest (Tshitoyan et al., 2019; Luca et al., 2022).
+    5.6 Stability with Respect to Anchor Selection Variability
+    The SPA routine chooses anchors that are highly used and orthogonal to those already selected.
+The analyst looks through the resulting anchors and the topics they define to assess their utility in measuring the pre-specified, hypothesized relationships.
+As a result the anchors selected by SPA (and used for our hypotheses) may vary if a new text sample is obtained.
+However this is different from the usual selection uncertainty problem which is cast as follows.
+Consider selecting an active subset of $x\subset X$ in regression $Y=f(X\beta)+\epsilon$.
+Here one might use LASSO or a stepwise routine which then requires post selection inference to produce interval estimates for the corresponding $\beta$ elements that properly account for the model selection.
+Our problem is inherrently different in that we are selecting appropriate values of $y\subset Y$ that suit our externally generated hypotheses.
+When performed using SPA, this action is taken without consideration of $X$.
+In text modelling all pre-processing steps impact the analytic outcome, whether it is converting words to lowercase, fixing typos, handling n-grams or stopwords.
+NMF decomposition and selection of anchors is a form of pre-processing step to be assessed for quality by the analyst taken before modelling begins.
+Analysis is then performed conditional on these decisions.
+Pre-processing is performed prior to considering covariate effects.
+    In the absence of a particular anchor word, an important topic will not vanish, but instead will evolve through a different anchor label and perhaps a different number of meaningful topics for the hypothesized relationships.
+In an extreme demonstration of the impact of changing anchors, we reconsider the problem of assessing a regional effect on hoppiness (H11 - H13) from Section 5.3, while forcing SPA to exclude {grapefruit, tropical, pine} from being considered as anchors.
+As before, the analyst must examine the resulting topics and assess their validity towards the inferential goal.
+In this experiment, we find the replacement hoppiness topics to be less focused.
+The important terms within selected anchors {citrus, fruit, orange, piney} are listed in table reftable:50topicexclusions.
+As before we find no significant deviation from the British Columbia baseline at the 5% level.
+A more thorough exploration of any potential impact of pre-processing is left for future work.
+    topic defining terms
+    citrus pine hoppy smooth rind appearance grapefruit peel
+    fruit tropical juicy mango apricot peach passion guava
+    orange grapefruit pine citrusy lingering peel crisp oily
+    piney pine grapefruit hoppy resin oily smooth copper
+    Table 6: Topics used to re-test regional differences in hoppyness discourse while excluding the selection of anchors grapefruit, tropical, pine.
+    6 Discussion
     In natural sciences, measurements are selected based on available equipment and often hypotheses must be assessed using observable proxies of interest.
@@
 Data curation and cleaning decisions require care and will impact results.
-When exploring regional differences in beers, all mentions of geography had to be removed from the text to prevent mention of Ontario in the text as being the driver for the Ontario covariate effect.
-With text especially, data decisions impact the outcome similar to how tuning experimental equipment in lab sciences must be conditioned upon when analyzing data.
-    The selection of topics as dependent variables requires care in how hypotheses are crafted to avoid the perils of using the same dataset for hypothesis generation and testing (Egami et al., 2022).
+When exploring regional differences in beers, all mentions of geography had to be removed from the text to prevent mention of Ontario or geographic proxies as being the driver for the Ontario covariate effect.
+Data cleaning decisions impact the outcome similar to how tuning experimental equipment in lab sciences must be conditioned upon when analyzing data.
+    The selection of topics requires care in how hypotheses are crafted to avoid the perils of using the same dataset for hypothesis generation and testing (Egami et al., 2022).
 The Beer hypotheses were constructed from external insights and interests around flavours, however determining anchor based topics that measure attributes of interest requires examination of $\Phi$, which in turn is based on the data.
-Reproducing the tests on a new independent dataset could be performed by bringing $\Phi$ or by carrying forward anchors.
-Even then, terminological drift suggests that application of the topic metric requires consideration of the utility of $\Phi$ within the new scientific context.
+The simulation study of section 4.2 suggests that using $\Phi$ from the NMF decomposition of an external independent corpus will increase the MSE relative to simply carrying forward the anchor terms and applying the decomposition on the new dataset.
+We expect that this will depend on the distance between word probabilities across the datasets though a thorough exploration thereof and it’s implications for experimental design is beyond the scope of this work.
+    While this paper considers OLS and Beta regression models, more complex regression models could be used.
+Generalized Additive Models or Deep Learning for Beta distributed data are natural extensions that follow from the BRETT framework.
     Increased discourse about a flavour topic may or may not represent actual difference in flavour probability.
@@
 More recent advances in Large Language Models have shown promise in categorizing text, however PTMs provide a magnitude of an observable feature rather than a binary categorical presence or absence.
-However there is considerable potential for exploiting LLMs as part of the data pre-processing and cleaning pipelines.
-    The main advantage to the BRETT approach is its stability.
-The NMF decomposition with anchor words is not subject to random seeds.
-Anchor words can be found automatically or (partially) provided by the analyst to fine tune hypotheses.
-The stability in BRETT allows covariate effects to be directly modelled and uses fewer approximations than STMs.
+There remains considerable potential for exploiting LLMs as part of the data pre-processing and cleaning pipelines.
+Embedding models such as BERT define lower dimensional latent word spaces that map words closer or further apart depending on their semantic meaning.
+As the vocabulary becomes more general, these tools are expected to provide considerable advantages in text modelling.
+    In many text applications time is an important factor, and to properly consider serial correlation in discourse, this should be considered within the NMF decomposition.
+Non-Negative Tensor Factorization may provide a useful starting point for future work on time-topic decomposition of text.
     Appendix A PTM and STM topics